### v1

# Permissible Knowledge Pooling

Abstract
Information pooling has been extensively formalised across various logical frameworks in distributed systems [1, 23, 11], characterized by diverse information-sharing patterns. These approaches generally adopt an intersection perspective, aggregating all possible information, regardless of whether it is known or unknown to the agents. In contrast, this work adopts a unique stance, emphasising that sharing knowledge means distributing what is known, rather than what remains uncertain. This paper introduces a dynamic logic for knowledge pooling or sharing and further discusses a potential framework for permissible knowledge pooling.
keywords:
Information Pooling, Distributed Knowledge, Knowledge Pooling, Permissible Knowledge Sharing.
1 Motivation
What disclosures are considered sensitive information leaks? What messages can be shared with a cooperative partner without breaching confidentiality agreements? In the processes of compliance checking, legal decision-making, and responsibility allocation, it is necessary not only to grasp the ethical ramifications but also to understand the types of epistemic information at play and their interaction with ethical considerations.
A notable gap exists in our understanding of the nature of epistemic information produced by these ethically questionable actions and the manner in which this information can shed light on the suitability or unsuitability of such actions. To address the above questions, existing research tends to focus either on the mechanics of information transfer inherent in such communicative acts [26, 6, 20, 27], or on the ethical considerations presumed by such actions [10, 9]. However, these approaches often overlook a comprehensive perspective.
The mishandling of sensitive information [3], especially in the context of multi-agent interactions, exemplifies the complexities involved in addressing such issues. This challenge underscores the need for a more integrated approach that considers both the mechanics of information exchange and the ethical implications thereof.
Example 1.1 (Sensitive Information).
A customer purchases services from a company to analyse user behaviors on websites. The company offers two servers, $a$ and $b$, each providing different type of information. Server $a$ tracks how users ($p$) are associated with user IDs ($q$), while server details how these user IDs correspond to the websites they visit ($r$). While the company is able to sell access to both servers, GDPR regulations prohibit selling both simultaneously, as doing so would result in a breach of privacy by revealing users’ personal website visitation patterns to the customer.
The knowledge provided by the servers can be represented as follows:
  • $K_{a}(p\to q)$ and
  • $K_{b}(q\to r)$.
GDPR restricts the customer from acquiring knowledge that links users directly to the websites they visit:
  • $K_{c}(p\to q)$.
Given that the business transaction requires at least one server to relay information to the customer, how can we construct a model of permissible knowledge transfer to investigate the underlying mechanisms?
Research Question: How do we determine the ethical considerations in sharing information?
This research question unfolds into two distinct inquiries:
[…]
    What mechanisms underpin the act of sharing information?
[…]
    How do the epistemic and ethical implications of communication converge to define the permissibility of such actions?
This work will address these questions as outlined below. Section 2 will introduce a dynamic model of knowledge sharing, highlighting the role of interpresonal dependence in the exchange of information. Following the tradition of dynamic deontic logic [24], Section 3 will develop a new framework to differentiate between the permissions to know and to share information. Section 4 will review relevant literature and Section 5 will provide concluding remarks.
2 Core Theory
2.1 Static Information
We adopt the standard approach of distributed knowledge [26] to examine the static dimension of information distribution and sharing. Within this framework, the modality $K_{a}$ represents the individual knowledge of agent $a$, while the modality $D_{G}$ signifies the distributed knowledge within a group $G$. This allows us to express that “Agent $a$ knows $\varphi$” using $K_{a}\varphi$ and that “Group $G$ (distributively) knows $\varphi$ if every members shares their knowledge” with $D_{G}\varphi$.
Definition 2.1 (Language).
The language $\mathcal{L}$ of knowledge is defined as follows:
$\varphi:=p\mid\neg\varphi\mid\varphi\wedge\varphi\mid K_{a}\varphi\mid D_{G}\varphi$
where $p\in Prop$ is an element of the (countable) set $Prop$ of atomic propositions, $a\in\mathcal{I}$ is an element of the (finite) set $\mathcal{I}$ of agents, and $G\subseteq\mathcal{I}$ is a subset of agents.
As usual, the dual modality $\hat{K_{a}}\varphi$ of individual knowledge is defined as $\neg K_{a}\neg\varphi$, which express the possibility of knowing individually. The “everybody knows” knowledge for group $G$, denoted as $E_{G}\varphi$, can be defined as $\bigwedge_{a\in G}K_{a}\varphi$.
Definition 2.2 (Models).
A structure $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},V\rangle$ is a model when:
  • $W$ is a non-empty set of possible states;
  • $R_{a}\subseteq W\times W$ is an equivalent relation;
  • $V:Prop\to 2^{W}$ is a valuation function.
The truth conditions of Boolean formulas and knowledge modalities are defined as usual [26]:
[…]
where $D_{G}=\bigcap_{a\in G}R_{a}$, $R_{a}[w]=\{u\in W\mid(w,u)\in R_{a}\}$, and $||\varphi||_{M}=\{w\in W\mid M,w\models\varphi\}$. When the model $M$ is clear from the context, we simplify the notation $||\varphi||_{M}$ to $||\varphi||$.
The axiomatization of these static modalities of knowledge is standard. The individual modality $K_{a}$ is a S5 modality and the modality of distributed knowledge characterized by the KT5 axiomatic framework [7].
Example 2.3 (Sensitive Knowledge, Cont’d).
The model $M$,as shown in Figure (1a), demonstrates the initial knowledge states of two servers $a$ and $b$, alongside the customer $c$. The expression $K_{a}(p\to q)$ clarifies that “Server $a$ know how users correlate with user IDs”. This is true at all states except for $w_{3}$ and $w_{4}$. Similarly, the expression $K_{b}(q\to r)$, which states “Server $b$ knows how user IDs correlate with websites”, is true at every state except for $w_{7}$ and $w_{8}$. The customer $c$, however, lacks solid knowledge regarding the mapping from $p$ to $q$ and from $q$ to $r$, which is illustrated in Figure (1b).
Upon the sharing of knowledge between servers $a$ and $b$, this act of pooling information results in the distributed knowledge regarding the mapping from $p$ users to $r$ websites. Specifically, the distributed knowledge $D_{\{a,b\}}(p\to r)$ holds true exclusively at states $w_{1},w_{2},w_{5}$, and $w_{6}$, illustrated in Figure (1a).
$w_{7}$$p,q,\neg r$$w_{8}$$\neg p,q,\neg r$$w_{2}$$\neg p,q,r$$w_{1}$$p,q,r$$w_{5}$$\neg p,\neg q,\neg r$$w_{6}$$\neg p,\neg q,r$$w_{4}$$p,\neg q,\neg r$$w_{3}$$p,\neg q,r$$bc$$bc$$ac$$ac$$bc$$ac$$ac$$bc$$abc$$abc$$abc$$abc$
(a) The initial knowledge of agents $abc$
$w_{7}$$p,q,\neg r$$w_{8}$$\neg p,q,\neg r$$w_{2}$$\neg p,q,r$$w_{1}$$p,q,r$$w_{5}$$\neg p,\neg q,\neg r$$w_{6}$$\neg p,\neg q,r$$w_{4}$$p,\neg q,\neg r$$w_{3}$$p,\neg q,r$$c$$c$$c$$c$$c$$c$$c$$c$$c$$c$$c$$c$
(b) The initial knowledge of agent $c$
Figure 1: The initial state of knowledge: $K_{a}(p\to q)$ and $K_{b}(q\to r)$.
2.2 Knowledge Sharing: A New Type of Information Pooling
Information flows from one agent to another, called information pooling, typically involves aggregating all available information from the participating agents through a process of information intersection [1, 23, 11], without necessarily confirming whether the information constitutes knowledge. To ensure the accuracy and integrity of communication, this section posits that information pooling ought to involve the sharing of knowledge. To express the concept of knowlege sharing between two agents, where agent $b$ receives agent $a$’s knowledge, we introduce the notation $a\smalltriangleright b$. Thus, the symbol $[a\smalltriangleright b]$ represents the act of sender $a$ conveying its knowledge to receiver $b$. Now we add the formula $[a\smalltriangleright b]\varphi$ into the language $\mathcal{L}$ and extend it into $\mathcal{L}^{+}$, and interpret this formula as follows:
Definition 2.4 (Updated Model).
Let $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},V\rangle$ be a model and $w\in W$ be a point in $M$. Then:
$M,w\models[a\smalltriangleright b]\varphi\textrm{ iff }M|_{a\smalltriangleright b},w\models\varphi$
where $M|_{a\smalltriangleright b}=\langle W,\{R^{*}_{i}\}_{i\in\mathcal{I}},V\rangle$ is a $(a\smalltriangleright b)$-updated model of knowledge sharing such that
[…]
    If $i\neq b$, then $R^{*}_{i}=R_{i}$;
[…]
    If $i=b$, then
    $uR^{*}_{b}s\textrm{ iff }uR_{b}s\textrm{ and }\forall p\in K_{a}u\,(u\in||p||_{M}\Leftrightarrow s\in||p||_{M})$
    where $K_{a}u=\{p\in Prop\mid M,u\models K_{a}p\}$.
Our updated mechanism is directional: It only eliminates those epistemic relations accessible to the recipient that do not align with the sender’s knowledge. In addition, this updating process is syntactic in nature: The model undergoes modification based on expressions that represent the sender’s knowledge.
The update process of knowledge sharing implies two key points. Firstly, it categorizes the recipient’s knowledge based on the knowledge provided by the sender, as shown in Lemma 2.5.2. Secondly, this update does not change the truth values of Boolean propositions, a principle demonstrated in Lemma 2.5.3.
Lemma 2.5.
Let $M|_{a\smalltriangleright b}=\langle W,\{R^{*}_{i}\}_{i\in\mathcal{I}},V\rangle$ be a $(a\smalltriangleright b)$-updated model.
[…]
    $M|_{a\smalltriangleright b}$ is still a model.
[…]
    Let $Prop^{+}$ be the set of Boolean propositions that only construct from atomic propositions, negation, and conjunction.
    $uR^{*}_{b}s\textrm{ iff }uR_{b}s\textrm{ and }\forall\varphi\in K_{a}^{+}u\,(u\in||\varphi||_{M}\Leftrightarrow s\in||\varphi||_{M})$
    where $K_{a}^{+}u=\{\varphi\in Prop^{+}\mid M,u\models K_{a}\varphi\}$.
[…]
    Let $u\in W$. So, $M,u\models\varphi$ iff $M|_{a\smalltriangleright b},u\models\varphi$, where $\varphi\in Prop^{+}$.
[…]
    $R_{a}\cap R_{b}\subseteq R^{*}_{b}$.
[…]
    $D_{G}=D_{G}^{*}$.
The properties of knowledge pooling allow us to simplify expressions that involve the updated modalities $[a\smalltriangleright b]$. Specifically, axioms A, C, N, $U_{i}$ ($1\leq i\leq 5$) enable the removal of knowledge pooling modalities when they interact with other boolean operators and knowledge modalities. Notice that the application of axiom $U_{1}$ is restricted to Boolean propositions.
Proposition 2.6.
Here are some validities about the update of $a\smalltriangleright b$:
[…]
    A: $[a\smalltriangleright b]p\leftrightarrow p$
[…]
    C: $[a\smalltriangleright b](\varphi\wedge\psi)\leftrightarrow[a\smalltriangleright b]\varphi\wedge[a\smalltriangleright b]\psi$, where $\varphi,\psi\in\mathcal{L}$
[…]
    N: $[a\smalltriangleright b]\neg\varphi\leftrightarrow\neg[a\smalltriangleright b]\varphi$ for any $\varphi\in\mathcal{L}$
[…]
    $U_{1}$: $K_{a}\varphi\to[a\smalltriangleright b]K_{b}\varphi$, where $\varphi\in Prop^{+}$
[…]
    $U_{2}$: $K_{b}\varphi\to[a\smalltriangleright b]K_{b}\varphi$ for $\varphi\in\mathcal{L}$
[…]
    $U_{3}$: $K_{c}\varphi\to[a\smalltriangleright b]K_{c}\varphi$ where $\varphi\in\mathcal{L}$ and $c\neq b$
[…]
    $U_{4}$: $[a\smalltriangleright b]K_{b}\varphi\to D_{\{a,b\}}[a\smalltriangleright b]\varphi$ where $\varphi\in\mathcal{L}$
[…]
    $U_{5}$: $[a\smalltriangleright b]D_{G}\varphi\leftrightarrow D_{G}[a\smalltriangleright b]\varphi$ where $\varphi\in\mathcal{L}$
[…]
    K-axiom: $[a\smalltriangleright b](\varphi\to\psi)\to([a\smalltriangleright b]\varphi\to[a\smalltriangleright b]\psi)$ where $\varphi,\psi\in\mathcal{L}$
[…]
    NEC Rule: If $\vdash\varphi$ then $\vdash[a\smalltriangleright b]\varphi$ where $\varphi\in\mathcal{L}$
Lemma 2.5.3 highlights that the process of knowledge pooling preserves the truth values of Boolean propositions. Here, furthermore, axioms A, C, and N explore a broader concept: The update process inherent in knowledge pooling maintains the logical behavior of boolean operators over the updated modalities.
Axioms $U_{1}$ to $U_{3}$ explain the process through which the knowledge of agent $a$ is accurately conveyed and integrated into the knowledge of another agent. Thus, the knowledge pooling from agent $a$ to agent $b$ augments the knowledge of the recipient $b$, without similar influence on the sender $a$ nor any third party such as agent $c$. This represents a unidirectional flow of information. Proposition 2.8.2 further shows how information can be flowed from one agent to another via the pooling process.
Axioms $U_{4}$ and $U_{5}$ indicate that the knowledge derived from the pooling process contributes to the collective understanding shared from the sender to the recipient, as shown by $U_{4}$. However, it is not implied that all distributed knowledge can be generated from the knowledge pooling process. This distinction is clarified in Example 2.7 presented below. Through this, knowledge pooling reaches a new type of collective knowledge that bridges individual knowledge and distributed knowledge.
Example 2.7.
Figure 2 illustrates a model $M=\langle W,R_{a},R_{b},V\rangle$ representing the individual knowledge of agents $a$ and $b$, where both agents only know $p$:
  • $W=\{s_{0},s_{1},s_{2},s_{3}\}$,
  • $R_{a}=\{(s,s)\mid s\in W\}\cup\{(s_{0},s_{1}),(s_{1},s_{0})\}$,
  • $R_{b}=\{(s,s)\mid s\in W\}\cup\{(s_{0},s_{2}),(s_{0},s_{3}),(s_{2},s_{0}),(s_{0},s_{3})\}$,
  • $V(p)=W$, $V(q)=\{s_{0},s_{3}\}$, and $V(r)=\{s_{0},s_{2}\}$.
Following the process of knowledge pooling, whether the direction is from from $a$ to $b$ or vice versa, it is not possible to generate any distributed knowledge beyond $p$ between these two agents. Essentially, the agents only share what they have known (e.g. $p$) and do not extend to what they are uncertain (e.g. $q,r$):
  • $M,s_{0}\not\models[a\smalltriangleright b]K_{b}(q\wedge r)$;
  • $M,s_{0}\not\models[b\smalltriangleright a]K_{a}(q\wedge r)$;
  • $M,s_{0}\models D_{\{a,b\}}(q\wedge r)$.
Distributed knowledge marks a boundary case of information pooling, characterized by the aggregation of all conceivable information, even if those are uncertain. Conversely, knowledge pooling is specifically aimed at gathering established or known information from the others.
$s_{0}$ $p,q,r$$s_{1}$ $p,\neg q,\neg r$$s_{2}$ $p,\neg q,r$$s_{3}$ $p,q,\neg r$$a$$b$$b$
Figure 2: Knowledge sharing: Beyond the aggregation of arbitrary information for distributed knowledge.
The unidirectional flow of knowledge pooling can be transferred through multi-agent interactions.
Proposition 2.8.
The following validities clarify the transitions of knowledge pooling from one agent towards the third party:
[…]
    $[a\smalltriangleright b][b\smalltriangleright c]K_{c}\varphi\to[a\smalltriangleright c]K_{c}\varphi$ where $\varphi\in Prop^{+}$
[…]
    $[a\smalltriangleright b]K_{b}\varphi\to[a\smalltriangleright b][b\smalltriangleright c]K_{c}\varphi$, where $\varphi\in Prop^{+}$
In this multi-agent context, our goal is to illustrate that the resolution of knowledge within the knowledge pooling process is not merely a matter of aggregating information arbitrarily.
Proposition 2.9 (A New Type of Resolution: Knowledge Resolution).
Consider a model $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},V\rangle$. The information resolution operator $R^{i}_{G}\varphi$, which represents the pooling of arbitrary information, is defined as per existing literature [1]:
$M,w\models R^{i}_{G}\varphi\textrm{ iff }M|_{G},w\models\varphi,$
where $M|_{G}=\langle W,\{R_{a}|_{G}\}_{a\in\mathcal{I}},V\rangle$ such that
$R_{a}|_{G}=\begin{cases}\bigcap_{b\in G}R_{b}&\quad a\in G,\\ R_{a}&\quad\textrm{otherwise}.\end{cases}$
While, we can define the notion of knowledge resolution by our operator of knowledge pooling:
$R_{G}\varphi:=[a\smalltriangleright b_{1}]\cdots[b_{n-1}\smalltriangleright b_{n}][b_{n}\smalltriangleright b_{n-1}]\cdots[b_{1}\smalltriangleright a]\varphi$
where $G=\{a,b_{1},\dots,b_{n}\}$.
Now we have these results:
[…]
    $K_{a}\varphi\to R_{G}E_{G}\varphi$, where $\varphi\in Prop^{+}$ and $G=\{a,b_{1},\dots,b_{n}\}$.
[…]
    $R_{G}E_{G}\varphi\to R^{i}_{G}E_{G}\varphi\leftrightarrow D_{G}R^{i}_{G}\varphi$, where $\varphi\in Prop^{+}$ and $a\in G$.
Example 2.10.
The model $M$ in Figure 2 is a counter-example of $R^{i}_{G}E_{G}\varphi\to R_{G}E_{G}\varphi$. In this model, the information resolution pools all consistent but unknown information gathering from agent $a$ and $b$ to have $R_{a}\cap R_{b}=\{(s_{0},s_{0})\}$. Therefore,
  • $M,s_{0}\models R^{i}_{\{a,b\}}E_{\{a,b\}}(p\wedge q\wedge r)$.
However, the knowledge resolution results
$R^{*}_{b}=\{(s_{0},s_{2}),(s_{0},s_{3})\}\cup\{(s,s)\mid s\in W\}.$
This leads to
  • $M,s_{0}\models R_{\{a,b\}}E_{\{a,b\}}p$,
  • $M,s_{0}\not\models R_{\{a,b\}}E_{\{a,b\}}q$, and
  • $M,s_{0}\not\models R_{\{a,b\}}E_{\{a,b\}}r$.
This Example 2.10 demonstrates that knowledge pooling results in a distinct form of collective knowledge, diverging from the distributed knowledge generated through information resolution. The relationship among individual knowledge, the knowledge achieved by knowledge pooling, and the knowledge obtained by information pooling can be outlines as follows:
$K_{a}\varphi\to[a\smalltriangleright b]K_{a}\varphi\to R_{G}E_{G}\varphi\to R^{i}_{G}E_{G}\varphi\leftrightarrow D_{G}R^{i}_{G}\varphi$
where $\varphi\in Prop^{+}$.
An important aspect of knowledge pooling is that, despite the various directions in which it can occur, there always exists a resolution that obtains all knowledge from the participating agents, as illustrated in Example 2.11 below.
Example 2.11 (Sensitive Information, Cont’d).
Starting with the initial model presented in Figure 1, knowledge sharing towards the customer $c$ can proceed in two directions: either server $a$ or server $b$ initiates sharing their knowledge. This leads to distinct updated models, as shown in Figure 3:
  • $M,w_{1}\models[a\smalltriangleright c]K_{c}(p\to q)$;
  • $M,w_{1}\models[b\smalltriangleright c]K_{c}(q\to r)$.
$w_{7}$$p,q,\neg r$$w_{8}$$\neg p,q,\neg r$$w_{2}$$\neg p,q,r$$w_{1}$$p,q,r$$w_{5}$$\neg p,\neg q,\neg r$$w_{6}$$\neg p,\neg q,r$$w_{4}$$p,\neg q,\neg r$$w_{3}$$p,\neg q,r$$c$$c$$c$$c$$c$$c$$c$$c$
(a) The impact of $a$’s knowledge sharing on $c$’s knowledge.
$w_{7}$$p,q,\neg r$$w_{8}$$\neg p,q,\neg r$$w_{2}$$\neg p,q,r$$w_{1}$$p,q,r$$w_{5}$$\neg p,\neg q,\neg r$$w_{6}$$\neg p,\neg q,r$$w_{4}$$p,\neg q,\neg r$$w_{3}$$p,\neg q,r$$c$$c$$c$$c$$c$$c$$c$$c$
(b) The impact of $b$’s knowledge sharing on $c$’s knowledge.
Figure 3: Two different directions of knowledge sharing
Within the framework of knowledge resolution, the concept of “everybody knows” knowledge [26] is achievable, as shown in Figure 4:
  • $M,w_{1}\models[a\smalltriangleright b][b\smalltriangleright c]E_{\{a,b,c\}}(p\to q)$;
  • $M,w_{1}\not\models[a\smalltriangleright b][b\smalltriangleright c]E_{\{a,b,c\}}(q\to r)$;
  • $M,w_{1}\models[b\smalltriangleright a][a\smalltriangleright c]E_{\{a,b,c\}}(q\to r)$;
  • $M,w_{1}\not\models[b\smalltriangleright a][a\smalltriangleright c]E_{\{a,b,c\}}(p\to q)$;
  • $M,w_{1}\models R_{\{a,b,c\}}(E_{\{a,b,c\}}(p\to q)\wedge E_{\{a,b,c\}}(q\to r)\wedge E_{\{a,b,c\}}(p\to r))$.
$w_{7}$$p,q,\neg r$$w_{8}$$\neg p,q,\neg r$$w_{2}$$\neg p,q,r$$w_{1}$$p,q,r$$w_{5}$$\neg p,\neg q,\neg r$$w_{6}$$\neg p,\neg q,r$$w_{4}$$p,\neg q,\neg r$$w_{3}$$p,\neg q,r$$c$$c$$c$$c$
Figure 4: When agent $c$ has received all knowledge both from $a$ and $b$.
The concept of knowledge resolution highlights a crucial insight from knowledge pooling: its capacity to facilitate a consensus among a group  [^1] If we incorporate the notion of common knowledge [26] into our framework, we can validate the following statement about our knowledge resolution: $R_{G}E_{G}\varphi\leftrightarrow R_{G}C_{G}\varphi,$ where $C_{G}$ is the modality of common knowledge in group $G$.. However, a knowledge resolution may not always allowed, particularly in scenarios involving sensitive information or issues arising from epistemic implications. The next section will propose a formalization for permissible knowledge pooling, aimed at addressing the gap between epistemic and ethical implications.
3 Permissible Knowledge Pooling
3.1 Permission to Know: Static Norms
To capture ethical implications within our model, we incorporate the concept of ideal knowledge transition. In standard deontic logic, the Kanger reduction typically employs a propositional constant to identify states within the model as permissible or impermissible [14]. To examine the interplay between epistemic and ethical implications, we introduce a propositional constant that specifies which pairs of states in the model, referred to as knowledge transitions, are permissible. Those knowledge transitions considered permissible are thus classified as ideal knowledge transitions.
We expand our framework of knowledge pooling as described below. A structure $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},O,V\rangle$ is a deontic model when $M$ is a model in additional with a symmetric relation $O\subseteq W\times W$ with the property that $\emptyset\neq O\subseteq\bigcup_{a\in\mathcal{I}}R_{a}$. To represent an ideal transition, we introduce the constant ${\sf O}$ and extend our language $\mathcal{L}$ to $\mathcal{L}^{O}$, indicating the presence of a permissible way to know – suggesting that it is okay to gain knowledge in such a manner. This constant is interpreted at a point $w$ in a deontic model $M$ as follows.
$M,w\models{\sf O}\textrm{ iff }O[w]\neq\emptyset.$
This constant has this validity:
${\sf O}\to\hat{K_{a}}{\sf O}$
for some agent $a$ [^2] By incorporating the concept of common knowledge, the validity expressed by $C_{\mathcal{I}}\bigvee_{a\in\mathcal{I}}O$ captures the characteristics of $O\subseteq\bigcup_{a\in\mathcal{I}}R_{a}$ and $O\neq\emptyset$.
Example 3.1 (Sensitive Information, Cont’d).
Figure (1a) illustrates the state pairs $(w_{1},w_{3})$, $(w_{3},w_{5})$, $(w_{5},w_{7})$, $(w_{7},w_{1})$, $(w_{2},w_{4})$, $(w_{4},w_{6})$, $(w_{6},w_{8})$, $(w_{8},w_{2})$, along with their corresponding symmetric pairs, constitute the ideal relation $O$. These pairs are thus considered as permissible pathways of obtaining knowledge.
We introduce a static notion termed as weak permission, denoted by $P_{a}\varphi$, and define it as $K_{a}\varphi\wedge\hat{K_{a}}{\sf O}$. Thus, $P_{a}\varphi$ at a point $w$ in the deontic model $M$ is interpreted as:
$M,w\models P_{a}\varphi\textrm{ iff }R_{a}[w]\subseteq||\varphi||\textrm{ and }\exists u\in R_{a}[w]\,\textrm{ s.t. }O[u]\neq\emptyset$
This implies that when an agent $a$ is permitted to acquire knowledge, it means that there is at least one ideal way for her to access her knowledge. The formula $P_{a}\varphi$ is designed to captures the concept of having permission to know. It asserts that an agent $a$ is permitted to know $\varphi$ if and only if, $\varphi$ is part of agent $a$’s knowledge base, and there is at least one permissible way of knowing it. This concept of “permission to know” can be expressed using a static language.
Example 3.2 (Sensitive Information, Cont’d).
After server $a$ conveys her knowledge to customer $c$, it becomes permissible for the agent $c$ to know $p\to q$. Likewise, when server $b$ shares his knowledge to customer $c$, it becomes permissible for $c$ to know $q\to r$. These permissions are true at point $w_{1}$ in the models illustrated in Figure 3:
  • $M|_{a\smalltriangleright c},w_{1}\models P_{c}(p\to q)$;
  • $M|_{b\smalltriangleright c},w_{1}\models P_{c}(q\to r)$.
However, in the model in Figure 4, it is not permissible for $c$ to know $p\to r$.
  • For instance, $(M|_{a\smalltriangleright c})|_{b\smalltriangleright c},w_{1}\not\models P_{c}(p\to r)$.
The static concept of permission to know displays characteristics derived from the principles of free choice permission [15] and individual knowledge:
Proposition 3.3.
Here are some validities:
[…]
    RFC: $P_{a}\varphi\wedge P_{a}\psi\to P_{a}(\varphi\vee\psi)$
[…]
    MC: $P_{a}(\varphi\wedge\psi)\leftrightarrow P_{a}\varphi\wedge P_{a}\psi$
[…]
    K: $P_{a}(\varphi\to\psi)\to(P_{a}\varphi\to P_{a}\psi)$
[…]
    D: $\neg P_{a}\bot$
[…]
    T: $P_{a}\varphi\to\varphi$ for all $a\in\mathcal{I}$
[…]
    B: $\varphi\to P_{a}\hat{P_{a}}\varphi$ for some $a\in\mathcal{I}$
[…]
    4: $P_{a}\varphi\to P_{a}P_{a}\varphi$ for all $a\in\mathcal{I}$
[…]
    RE: If $\vdash\varphi\to\psi$ then $\vdash P_{a}\varphi\to P_{a}\psi$
[…]
    NEC: If $\vdash\varphi$ then $\vdash P_{a}\varphi$
However, this concept avoid the infamous free choice permission paradox [15] and the 5 axiom of knowledge:
Example 3.4.
Here are some invalidities for all $a\in\mathcal{I}$:
[…]
    FCP1: $P_{a}(\varphi\vee\psi)\to P_{a}\varphi\wedge P_{a}\psi$
[…]
    FCP2: $P_{a}\varphi\to P_{a}(\varphi\wedge\psi)$
[…]
    5: $\hat{P_{a}}\varphi\to P_{a}\hat{P_{a}}\varphi$
Significantly, the permission to access knowledge represents a novel bridge linking individual knowledge with distributed knowledge, as described below:
$D_{G}(\varphi\wedge{\sf O})\to P_{a}\varphi\to K_{a}\varphi\to E_{G}\varphi$
The dual concept of “permission to know” – “ought to know” can be defined in the usual manner: $O_{a}\varphi:=\hat{P_{a}}\varphi$. Thus, “ought to know” $O_{a}\varphi$ corresponds to a conditional $(K_{a}\neg\varphi\to K_{a}\neg{\sf O})$. This interpretation of “ought to know” avoids the infamous Åqvist’s paradox.
3.2 Permission to Knowledge Pooling: Dynamic Norms
Now we define a new type of permission – the permission for one to communicate their knowledge towards another. This introduces a novel dynamic dimension to the concept of permission, particularly in terms of communication, that is represented as $P(a\smalltriangleright b)$. It denotes that agent $a$ is permitted to share their knowledge with that of agent $b$. More precisely, $P(a\smalltriangleright b)$ is defined as
$K_{a}\varphi\to[a\smalltriangleright b]P_{b}\varphi.$
Agent $a$ is allowed to transfer knowledge to agent $b$, if and only if, upon agent $a$ being allowed to possess knowledge of $\varphi$, the subsequent communication from $a$ to $b$ results in receiver $b$ is allowed to possess knowledge of $\varphi$. Due to each agent has its own constant of violation, $P(a\smalltriangleright b)$ may be false.
The extension of language $\mathcal{L}$ to include $P(a\smalltriangleright b)$ is denoted as $\mathcal{L}^{P}$, which is expanded to accommodate this dynamic type of permission.
Proposition 3.5.
The language $\mathcal{L}^{+}$ cannot be expressed by the language $\mathcal{L}^{P}$, and vice versa.
Example 3.6 (Sensitive Information, Cont’d).
Within the use of the language $\mathcal{L}^{P}$, we can further explore the case of sensitive information in Example 3.2. In the model $M|_{a\smalltriangleright c}$, server $b$ knows $q\to r$. When $b$ shares this knowledge with customer $c$, $c$ know it as well. However, there is no legitimate means for her to have obtained this knowledge.
  • $M|_{a\smalltriangleright c},w_{1}\models K_{b}(q\to r)$
  • $M|_{a\smalltriangleright c},w_{1}\not\models[b\smalltriangleright c]P_{c}(q\to r)$
From these observations, we deduce that the transfer of knowledge from $b$ to $c$ is impermissible.
  • $M|_{a\smalltriangleright c},w_{1}\not\models P(b\smalltriangleright c)$.
Knowledge pooling shares significant ties with the concept of ’‘permissions to share knowledge”, particularly within multi-agent interaction scenarios.
Proposition 3.7.
Here are some validities:
[…]
    $[a\smalltriangleright b][b\smalltriangleright c]\hat{K_{c}}{\sf O}\to[a\smalltriangleright c]\hat{K_{c}}{\sf O}$
[…]
    $P(a\smalltriangleright b)\wedge[a\smalltriangleright b]P(b\smalltriangleright c)\to P(a\smalltriangleright c)$
[…]
    If $\vdash K_{a}\varphi\leftrightarrow K_{b}\varphi$ then $P(a\smalltriangleright c)\leftrightarrow P(b\smalltriangleright c)$
Proposition 3.7.1 demonstrates that if it is permissible for agent $c$ to acquire knowledge through sequential sharing from $a$ to $b$ and then from $b$ to $c$, this permissibility remains even when the knowledge is directly pooled from $a$ to $c$. Proposition 3.7.2 explores a different facet of permissible knowledge pooling: if the transfer of knowledge is permissible from $a$ to $b$ and, after this transfer, similarly from $b$ to $c$, then a direct transfer from $a$ to $c$ is also permissible. Proposition 3.7.3 addresses the condition for replacing participants in permissible pooling: if the originating knowledge of the senders is identical, then the permissibility status of their knowledge pooling is the same. Therefore, Proposition 3.7.3 explains one scenario in which epistemic implications can affect ethical considerations. Example 3.8 below shows that substituting recipients with identical knowledge does not guarantee the preservation of the original permissibility status.
Example 3.8.
Here are some invalidities:
[…]
    If $K_{b}\varphi\leftrightarrow K_{c}\varphi$ then $P(a\smalltriangleright b)\leftrightarrow P(a\smalltriangleright c)$.
4 Related Work
4.1 Methodology of Link Deletion
Our model update, focused on $(a\smalltriangleright b)$-knowledge sharing, selectively eliminates all $b$-links within the model that do not align with agent $a$’s knowledge. In the field of model update and modification, deletion is widely recognized as a core strategy, as evidenced by various methodologies including public announcement [17, 18], semantic updates [28], product updates [5], upgrades [21], and sabotage updates [19] among others. Technical surveys detailing the results of their complexities can refereed to [26, 20, 2]. Link deletion, in particular, is essential for defining strategies in information pooling, where each method of update is designed for a specific function: Product updates [5, 26] are agent-oriented, upgrades [21] focus on formula-specific modifications, and sabotage methods [19, 20] target the removal of arbitrary links without regard to specific expression  [^3] A recent proposal introduces a language-specific approach to sabotage updates [2]. Furthermore, an exploration of updates from a language-specific perspective to those unrelated to language in sabotage modifications was presented in [22]. . In the domain of information pooling, algebraic intersection methods that consolidate all potential information relative to specific agents [1, 23, 11] stand out. Unlike previous approaches, our method remains agent-specific in the syntax in its deletion of links, but notably underscores the semantic role of knowledge formulas in the update process.
4.2 Social Dependence
Our knowledge pooling model brings to light a dimension of dependence within social interactions, a concept extensively examined in various formal theories. The notion of belief dependence, as introduced by Delgrande et al. [13], focuses on belief change and revision, thereby framing epistemic dependence. However, it is formalized within conditional logics employing a static language framework. On a different note, Boella et al. [8] have proposed a formal model designed to explain the power dynamics inherent in social dependence, rooted in an agent’s reliance on others to achieve their goals. This model of normative dependence utilises Castelfranchi’s methodology of dependence graphs [12], albeit without integrating a logical system.
In contrast, our framework for permissible knowledge pooling introduces a concept of dependence that not only underscores the dynamics of knowledge among individuals but also explores the capacity for maintaining the permissibility of information sharing between agents. This approach sheds light on the complex network of interpersonal dependencies, examined through both epistemic and normative lenses. Thus, this approach holds promise for future explorations into understanding social hierarchies and power dynamics.
4.3 Permission to Know
Our methodology for permissible knowledge pooling draws on the established logic of permission, as introduced by van der Meyden [24], which identifies permissibility across pairs of states concerning the authorization to acquire knowledge. Traditionally, most models have formalized the concept of permission to know by highlighting permissibility in relation to states. Moreover, inspired from Aucher et al. [3], several existing studies view permission to know as contingent upon permission to communicate—postulating that knowledge acquired through communication should not result in a breach of permissible states, as supported by the research of Balbiani et al. [4], Li et al.  [16], and Van Ditmarsch and Seban [25].
Our approach introduces a different perspective on the interplay between the permission to know and the permission to communicate, emphasising that the former precedes the latter. In essence, our framework suggests that we are granted the permission to know prior to receiving the authorization to communicate. This conceptualization of permissible knowledge pooling is distinctly defined through the lens of permission to know, offering a fresh viewpoint on its relationship with communicative permissions.
5 Conclusions
This study explores two key aspects of multi-agent communication: knowledge pooling and the permission to share knowledge. Unlike the conventional approach to information pooling, which aggregates arbitrary information to form collective knowledge, our framework selectively integrates only verified and known information during the pooling process. This approach introduces a novel perspective to information pooling, uncovering a unique form of collective knowledge that bridges the gap between individual and distributed knowledge.
Furthermore, our investigation offers a unique viewpoint on the interaction between epistemic and ethical considerations, successfully establishing a framework that explains how epistemic states together with our normative entitlement to knowledge influence our actions. By examining the relationship between the permission to know and the permission to share knowledge, we provide insightful observations on this interplay.
In our formal framework, we have investigated a range of validities and expressivities related to knowledge pooling, distributed knowledge, and the permission to know, emphasizing their roles within permissible knowledge pooling. We leave the exploration of axiomatization, computational complexity, and decidability to future technical research.
Our framework for permissible knowledge pooling introduces various challenges in the formalization of social reasoning within multi-agent interactions. For example, investigating the applicability of our concept of knowledge pooling – as a form of social dependence – to interpret social power is of significant interest. This includes examining the correlation between the extent of knowledge sharing and the structure of social hierarchies. Additionally, incorporating the concept of common knowledge into our framework presents an opportunity to explore deeper into the relationship between knowledge pooling and the achievement of social consensus.
Acknowledgement
This work is supported by the Fonds National de la Recherche of Luxembourg through the project Deontic Logic for Epistemic Rights (OPEN O20/14776480).
Appendix
Lemma .1.
Let $M|_{a\smalltriangleright b}=\langle W,\{R^{*}_{i}\}_{i\in\mathcal{I}},V\rangle$ be a $(a\smalltriangleright b)$-updated model.
[…]
    $M|_{a\smalltriangleright b}$ is still a model.
[…]
    Let $Prop^{+}$ be the set of Boolean propositions that only construct from atomic propositions, negation, and conjunction.
    $uR^{*}_{b}s\textrm{ iff }uR_{b}s\textrm{ and }\forall\varphi\in K_{a}^{+}u\,(u\in||\varphi||_{M}\Leftrightarrow s\in||\varphi||_{M})$
    where $K_{a}^{+}u=\{\varphi\in Prop^{+}\mid M,u\models K_{a}\varphi\}$.
[…]
    Let $u\in W$. So, $M,u\models\varphi$ iff $M|_{a\smalltriangleright b},u\models\varphi$, where $\varphi\in Prop^{+}$.
[…]
    $R_{a}\cap R_{b}\subseteq R^{*}_{b}$.
[…]
    $D_{G}=D_{G}^{*}$.
Proof .2.
[…]
    We only need to show that $R^{*}_{b}$ is still an equivalent relation.
    • Reflexivity: Due to the reflexivity of $R_{b}$ we have $uR_{b}u$ where any $u\in R_{b}[w]$. Due to Definition 2.4.2, it is obvious that the bi-direction holds. So $uR^{*}_{b}u$ for any $u\in W$.
    • Transitivity: Given $uR^{*}_{b}s$ and $sR^{*}_{b}v$, we know that $uR_{b}s$ and $sR_{b}v$. So we have $uR_{b}v$ because of the transitivity of $R_{b}$. According to the bi-direction in Definition 2.4.2, it is the case that $\forall p\in K_{a}u(u\in||p||_{M}\Leftrightarrow v\in||p||_{M})$. Now we have $uR^{*}_{b}v$.
    • Symmetry: Let $uR^{*}_{b}s$. So $uR_{b}s$. According to the bi-equivalence of Definition 2.4.2, it is not hard to imply $\forall p\in K_{a}s\,(s\in||p||_{M}\Leftrightarrow u\in||p||_{M})$. This obviously leads to $uR^{*}_{b}s$.
[…]
    We only need to show that from $uR^{*}_{b}s$ we can have the bi-direction for all formulas in $K_{a}^{+}w$. This can be proven by inducing on the complexity of formulas in $Prop^{+}$ in a standard way.
[…]
    Inducing on the complexity of $Prop^{+}$.
    • Basic case: When $\varphi=p\in Prop$, this bi-equivalence holds by the definition of the valuation function in $M|_{a\smalltriangleright b}$.
    • Induction hypothesis: Given a subformula $\psi\in Prop^{+}$ of $\varphi$, the bi-equivalence for $\psi$ holds.
    • Now the other cases such as either $\varphi=\neg\psi$ or $\varphi=\psi\wedge\psi^{\prime}$ can easily be proven due to the induction hypothesis in a standard way.
[…]
    This is obvious by Definition 2.4.2.
[…]
    This can be shown by the set-theoretic feature that both $D_{G}$ and $D^{*}_{G}$ are obtained by intersection.
Proposition .3.
Here are some validities about the update of $a\smalltriangleright b$:
[…]
    A: $[a\smalltriangleright b]p\leftrightarrow p$
[…]
    C: $[a\smalltriangleright b](\varphi\wedge\psi)\leftrightarrow[a\smalltriangleright b]\varphi\wedge[a\smalltriangleright b]\psi$, where $\varphi,\psi\in\mathcal{L}$
[…]
    N: $[a\smalltriangleright b]\neg\varphi\leftrightarrow\neg[a\smalltriangleright b]\varphi$ for any $\varphi\in\mathcal{L}$
[…]
    $U_{1}$: $K_{a}\varphi\to[a\smalltriangleright b]K_{b}\varphi$, where $\varphi\in Prop^{+}$
[…]
    $U_{2}$: $K_{b}\varphi\to[a\smalltriangleright b]K_{b}\varphi$ for $\varphi\in\mathcal{L}$
[…]
    $U_{3}$: $K_{c}\varphi\to[a\smalltriangleright b]K_{c}\varphi$ where $\varphi\in\mathcal{L}$ and $c\neq b$
[…]
    $U_{4}$: $[a\smalltriangleright b]K_{b}\varphi\to D_{\{a,b\}}[a\smalltriangleright b]\varphi$ where $\varphi\in\mathcal{L}$
[…]
    $U_{5}$: $[a\smalltriangleright b]D_{G}\varphi\leftrightarrow D_{G}[a\smalltriangleright b]\varphi$ where $\varphi\in\mathcal{L}$
[…]
    K-axiom: $[a\smalltriangleright b](\varphi\to\psi)\to([a\smalltriangleright b]\varphi\to[a\smalltriangleright b]\psi)$ where $\varphi,\psi\in\mathcal{L}$
[…]
    NEC Rule: If $\vdash\varphi$ then $\vdash[a\smalltriangleright b]\varphi$ where $\varphi\in\mathcal{L}$
Proof .4.
[…]
    The truth condition for atomic propositions is never changed, according to Definition 2.4.
[…]
    This can be proved by using Definition 2.4 many times.
[…]
    The negation about pooling behaves in this way:
[…]
    Let $M,w\models K_{a}\varphi$ and $wR^{*}_{b}u$ in the updated model $M|_{a\smalltriangleright b}$, where $\varphi\in Prop^{+}$. We need to show that $M|_{a\smalltriangleright b},u\models\varphi$. According to Lemma 2.5.2 and $wR^{*}_{b}u$, we know that $\forall\varphi\in K_{a}^{+}w\,(w\in||\varphi||_{M}\Leftrightarrow u\in||\varphi||_{M})$ (@). From $M,w\models K_{a}\varphi$ and T-axiom, we know that $M,w\models\varphi$. Then, by (@), it implies that $M,u\models\varphi$. Then, due to Lemma 2.5.3, we conclude that $M|_{a\smalltriangleright b},u\models\varphi$.
[…]
    This axiom is valid because $R_{b}^{*}\subseteq R_{b}$ by Definition 2.4.2.
[…]
    This axiom is valid because $R^{*}_{a}=R_{a}$ by Definition 2.4.1.
[…]
    This can be demonstrated by using Lemma 2.5.4.
[…]
    This can be shown by using Lemma 2.5.5.
[…]
    As Lemma 2.5.1 demonstrated, after a $(a\smalltriangleright b)$-update, the structure $M|_{a\smalltriangleright b}$ is still a model. It therefore concludes $\psi$ being true at $w$, when $\varphi\to\psi$ and $\varphi$ are both satisfied at the same point $w$ in the updated model $M|_{a\smalltriangleright b}$.
[…]
    If $\varphi$ is true everywhere in the model, then it simply ensures it is always true at both sides of the bi-direction in Definition 2.4.2 for every pair of worlds. This indicates that, after any $(a\smalltriangleright b)$-update, $\varphi$ is still true everywhere in the updated model.
Proposition .5.
The following validities clarify the transitions of knowledge pooling from one agent towards the third party:
[…]
    $[a\smalltriangleright b][b\smalltriangleright c]K_{c}\varphi\to[a\smalltriangleright c]K_{c}\varphi$ where $\varphi\in Prop^{+}$
[…]
    $[a\smalltriangleright b]K_{b}\varphi\to[a\smalltriangleright b][b\smalltriangleright c]K_{c}\varphi$, where $\varphi\in Prop^{+}$
Proof .6.
[…]
    First of all, in a $(a\smalltriangleright b)$-model $M$, it has $K^{+}_{b}w|_{a\smalltriangleright b}=Cn(K^{+}_{b}w\cap K^{+}_{a}w)$. Let $K^{+}_{b}w|_{a\smalltriangleright b}=K^{+}_{b}w^{*}$. Similarly, we have $K^{+}_{c}(w|_{a\smalltriangleright b})|_{b\smalltriangleright c}=Cn(K^{+}_{c}w\cap K^{+}_{b}w^{*})$ and $K^{+}_{c}w|_{a\smalltriangleright c}=Cn(K^{+}_{c}w\cap K^{+}_{a}w)$. So $K^{+}_{c}(w|_{a\smalltriangleright b})|_{b\smalltriangleright c}\subseteq K^{+}_{c}w|_{a\smalltriangleright c}$. Together with Lemma 2.5.3, this ensures the desired result.
[…]
    Let $M,w\models[a\smalltriangleright b]K_{b}\varphi$ where $\varphi\in Prop^{+}$. This is $M|_{a\smalltriangleright b},w\models K_{b}\varphi$. Because of Proposition 2.6.4, we know that $M|_{a\smalltriangleright b},w\models K_{b}\varphi\to[b\smalltriangleright c]K_{c}\varphi$. Now we conclude with $M|_{a\smalltriangleright b},w\models[b\smalltriangleright c]K_{c}\varphi$. This leads to the desired result $M,w\models[a\smalltriangleright b][b\smalltriangleright c]K_{c}\varphi$.
Proposition .7.
[…]
    $K_{a}\varphi\to R_{G}E_{G}\varphi$, where $\varphi\in Prop^{+}$ and $G=\{a,b_{1},\dots,b_{n}\}$.
[…]
    $R_{G}E_{G}\varphi\to R^{i}_{G}E_{G}\varphi\leftrightarrow D_{G}R^{i}_{G}\varphi$, where $\varphi\in Prop^{+}$ and $a\in G$.
Proof .8.
[…]
    We want to prove $K_{a}\varphi\to[a\smalltriangleright b_{1}]\cdots[b_{n-1}\smalltriangleright b_{n}][b_{n}\smalltriangleright b_{n-1}]\cdots[b_{1}\smalltriangleright a]E_{G}\varphi$. From Proposition 2.6, we have
    • $K_{a}\varphi\to[a\smalltriangleright b_{1}]K_{b_{1}}\varphi$
      (by Proposition 2.6.4)
    • $[b_{n}\smalltriangleright b_{n+1}]K_{b_{n+1}}\varphi\to[b_{n}\smalltriangleright b_{n+1}][b_{n+1}\smalltriangleright b_{n+2}]K_{b_{n+2}}\varphi$
      (by Proposition 2.8.2)
    • $[b_{n}\smalltriangleright b_{n+1}]K_{b_{n+1}}\varphi\to[b_{n}\smalltriangleright b_{n+1}][b_{n+1}\smalltriangleright b_{n+2}]K_{b_{n+1}}\varphi$
      (by Proposition 2.6.6, NEC, K, MP)
    And these validities together can lead to the desired result.
[…]
    First of all, when $b_{i}\in G$ ($1\leq i\leq n$)
    $R_{b_{i}}|_{G}=\bigcap_{b\in G}R_{b}\subseteq R^{*}_{b_{i}}$
    according to Definition 2.4. So the first implication holds. The second bi-direction also holds and its proof can refer to [1].
Proposition .9.
Here are some validities:
[…]
    $[a\smalltriangleright b][b\smalltriangleright c]\hat{K_{c}}{\sf O}\to[a\smalltriangleright c]\hat{K_{c}}{\sf O}$
[…]
    $P(a\smalltriangleright b)\wedge[a\smalltriangleright b]P(b\smalltriangleright c)\to P(a\smalltriangleright c)$
[…]
    If $\vdash K_{a}\varphi\leftrightarrow K_{b}\varphi$ then $P(a\smalltriangleright c)\leftrightarrow P(b\smalltriangleright c)$
Proof .10.
[…]
    As the proof for Proposition 2.8.1 shown, it is the case that $K_{c}(w|_{a\smalltriangleright b})|_{b\smalltriangleright c}\subseteq K_{c}w|_{a\smalltriangleright c}$. By Definition 2.4, it ensures that $((R_{c})|_{a\smalltriangleright b})|_{b\smalltriangleright c}[w]\subseteq(R_{c})|_{a\smalltriangleright c}[w]$, which leads us to the desired result.
[…]
    Let $M,w\models P(a\smalltriangleright b)$ and $M,w\models[a\smalltriangleright b]P(b\smalltriangleright c)$. This implies
    • If $M,w\models K_{a}\varphi$ then $M|_{a\smalltriangleright b},w\models P_{b}\varphi$ for any $\varphi$;
    • If $M,w\models[a\smalltriangleright b]K_{b}\varphi$ then $M,w\models[a\smalltriangleright b][b\smalltriangleright c]P_{c}\varphi$.
    So we have
    • If $M,w\models K_{a}\varphi$ then $M,w\models[a\smalltriangleright b][b\smalltriangleright c]P_{c}\varphi$ for any $\varphi$.
    From Proposition 2.8.1 we have:
    • $[a\smalltriangleright b][b\smalltriangleright c]K_{c}\varphi\to[a\smalltriangleright c]K_{c}\varphi$;
    While, from Proposition 3.7.1 we have
    • $[a\smalltriangleright b][b\smalltriangleright c]{\sf O}_{c}\to[a\smalltriangleright c]{\sf O}_{c}$.
    These two statements result
    • $[a\smalltriangleright b][b\smalltriangleright c]P_{c}\varphi\to[a\smalltriangleright c]P_{c}\varphi$.
    This leads to what we want to prove $M,w\models P(a\smalltriangleright c)$, which is
    • If $M,w\models K_{a}\varphi$ then $M,w\models[a\smalltriangleright c]P_{c}\varphi$ for any $\varphi$.