### v3

# Permissible Knowledge Pooling

Abstract
Information pooling has been extensively formalised across various logical frameworks in distributed systems [1, 30, 13, 8], characterized by diverse information-sharing patterns. These approaches generally adopt an intersection perspective, aggregating all possible information, regardless of whether it is known or unknown to the agents. In contrast, this work adopts a unique stance, emphasising that sharing knowledge means distributing what is known, rather than what remains uncertain. This paper introduces new modal logics for knowledge pooling and sharing, ranging from a novel language of knowledge pooling to a dynamic mechanism for knowledge sharing. It also outlines their axiomatizations and discusses a potential framework for permissible knowledge pooling.
keywords:
Information Pooling, Distributed Knowledge, Knowledge Pooling, Permissible Knowledge Sharing.
1 Motivation
What disclosures are considered sensitive information leaks? What messages can be shared with a cooperative partner without breaching confidentiality agreements? In the processes of compliance checking, legal decision-making, and responsibility allocation, it is necessary not only to grasp the ethical ramifications but also to understand the types of epistemic information at play and their interaction with ethical considerations.
A notable gap exists in our understanding of the nature of epistemic information produced by these ethically questionable actions and the manner in which this information can shed light on the suitability or unsuitability of such actions. To address the above questions, existing research tends to focus either on the mechanics of information transfer inherent in such communicative acts [33, 7, 26, 34], or on the ethical considerations presumed by such actions [12, 11]. However, these approaches often overlook a comprehensive perspective.
The mishandling of sensitive information [4], especially in the context of multi-agent interactions, exemplifies the complexities involved in addressing such issues. This challenge underscores the need for a more integrated approach that considers both the mechanics of information exchange and the ethical implications thereof.
Example 1.1 (Sensitive Information).
A customer purchases services from a company to analyse user behaviors on websites. The company offers two servers, $a$ and $b$, each providing different type of information. Server $a$ tracks how users ($p$) are associated with user IDs ($q$), while server $b$ details how these user IDs correspond to the websites they visit ($r$). While the company is able to sell access to both servers, GDPR regulations prohibit selling both simultaneously, as doing so would result in a breach of privacy by revealing users’ personal website visitation patterns to the customer.
The knowledge provided by the servers can be represented as follows:
  • $K_{a}(p\to q)$ and
  • $K_{b}(q\to r)$.
GDPR restricts the customer from acquiring knowledge that links users directly to the websites they visit:
  • $K_{c}(p\to r)$.
Given that the business transaction requires at least one server to relay information to the customer, how can we construct a model of permissible knowledge transfer to investigate the underlying mechanisms?
Research Question: How do we determine the ethical considerations in sharing information?
This research question unfolds into two distinct inquiries:
[…]
    What mechanisms underpin the act of sharing information?
[…]
    How do the epistemic and ethical implications of communication converge to define the permissibility of such actions?
This work will address these questions as outlined below. Section 2 will introduce a dynamic model of knowledge sharing, highlighting the role of interpresonal dependence in the exchange of information. Following the tradition of dynamic deontic logic [31], Section 3 will develop a new framework to differentiate between the permissions to know and to share information. Section 4 will discuss various results concerning expressivity, axiomatization, and completeness. Section 5 will review relevant literature and Section 6 will provide concluding remarks.
2 Theory of Knowledge
2.1 Static Information
We follow the standard approach of distributed knowledge [33] to examine the static dimension of information distribution and sharing. Within this framework, the modality $K_{a}$ represents the individual knowledge of agent $a$, while the modality $D_{G}$ signifies the distributed knowledge within a group $G$. This allows us to express that “Agent $a$ knows $\varphi$” using $K_{a}\varphi$ and that “Group $G$ (distributively) knows $\varphi$ if every members shares their knowledge” with $D_{G}\varphi$.
An agent’s knowledge can be augmented through the acquisition of insights from another. For instance, a customer’s understanding of user behaviours on websites is reinforced when one server shares its knowledge. This agent-dependent knowledge will be expressed by a new modality $K^{a}_{b}$. The expression $K^{a}_{b}\varphi$ represents that agent $b$ possesses knowledge of $\varphi$ dependent on the input from agent $a$.
Definition 2.1 (Language).
The language $\mathcal{L}$ of agent-dependent knowledge is defined as follows:
$\varphi:=p\mid\neg\varphi\mid\varphi\wedge\varphi\mid K_{a}\varphi\mid D_{G}\varphi\mid K^{a}_{b}\varphi$
where $p\in Prop$ is an element of the (countable) set $Prop$ of atomic propositions, $a\in\mathcal{I}$ is an element of the (finite) set $\mathcal{I}$ of agents, and $G\subseteq\mathcal{I}$ is a subset of agents.
As usual, the dual modality $\hat{K_{a}}\varphi$ of individual knowledge is defined as $\neg K_{a}\neg\varphi$, which expresses the possibility of knowing individually. The “everybody knows” knowledge for group $G$, denoted as $E_{G}\varphi$, can be defined as $\bigwedge_{a\in G}K_{a}\varphi$.
Definition 2.2 (Models).
A structure $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},V\rangle$ is a model when:
  • $W$ is a non-empty set of possible states;
  • $R_{a}\subseteq W\times W$ is an equivalence relation;
  • $V:Prop\to 2^{W}$ is a valuation function.
The truth conditions of Boolean formulas and knowledge modalities are defined as usual [33]:
[…]
where $D_{G}=\bigcap_{a\in G}R_{a}$, $R_{a}[w]=\{u\in W\mid(w,u)\in R_{a}\}$, and $||\varphi||_{M}=\{w\in W\mid M,w\models\varphi\}$. When the model $M$ is clear from the context, we simplify the notation $||\varphi||_{M}$ to $||\varphi||$. We define $\equiv_{a:w}\subseteq W\times W$ as an equivalence relation based on agent $a$’s basic knowledge at state $w$: $s\equiv_{a:w}u$ if and only if $s\in||\varphi||\Leftrightarrow u\in||\varphi||$ for all $\varphi\in K_{a}w$, where $K_{a}w=\{\varphi\in\mathcal{L}\mid M,w\models K_{a}\varphi\}$ is agent $a$’s knowledge at $w$. Now the agent-dependent knowledge is interpreted as follows:
$M,w\models K^{a}_{b}\varphi\text{ iff }(R_{b}\cap\equiv_{a:w})[w]\subseteq||\varphi||.$
The relation $R_{b}\cap\equiv_{a:w}$ is an equivalence relation.
The expression $K^{a}_{b}\varphi$ analyses $b$’s knowledge by integrating the relevant knowledge of $a$. This integration is interpreted by the intersection between $b$’s knowledge accessible relation $R_{b}$ and a corresponding categorization of agent $a$’s knowledge. Essentially, it can be understood as “Agent $b$ knows $\varphi$ when considering the knowledge from $a$ remain constant.” This approach of knowledge, grounded in equality, reflects the concept of ceteris paribus (with all other things else being equal) [27, 19, 18] in the perspective of epistemology.
These static modalities of knowledge are considered standard. The modality of individual knowledge $K_{a}$ and the modality of distributed knowledge are both characterized by the S5 axiomatic frameworks [9]. The agent-dependent modality $K^{a}_{b}$ is an S5 modality as well therefore, capturing the essence of knowledge. However, the interactions between them extend the standard S5 frameworks.
  • (Int) $K_{a}\varphi\to K^{b}_{a}\varphi$
  • (Cl) $K^{b}_{a}\varphi\to\bigvee_{\psi\in\mathcal{L}}(K_{a}\psi\wedge K_{b}(\psi\to\varphi))$
Details about axiomatizations of these modalities are presented in Section 4.2.
The following logical implication captures the relationship among these three types of knowledge:
$K_{b}\varphi\to K^{a}_{b}\varphi\to D_{\{a,b\}}\varphi.$
Agent-dependent knowledge serves as a bridge from individual to distributed knowledge, filling the connection from individual towards distributed knowledge. The two examples below further clarify this relationship.
Example 2.3 (Sensitive Information, Cont’d).
Figure 1 illustrates a model $M=\langle W,R_{a},R_{b},R_{c},V\rangle$ representing the individual knowledge of agents $a$ and $b$, where both agents only know $p$:
  • $W=\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$,
  • $R_{a}=\{(s,s)\mid s\in W\}\cup\{(s_{0},s_{1}),(s_{1},s_{0}),(s_{0},s_{2}),(s_{2},s_{0})\}$,
  • $R_{b}=\{(s,s)\mid s\in W\}\cup\{(s_{0},s_{3}),(s_{3},s_{0}),(s_{0},s_{4}),(s_{4},s_{0})\}$,
  • $R_{c}=R_{a}\cup R_{b}$
  • $V(p)=\{s_{0},s_{1},s_{3}\}$, $V(q)=\{s_{0},s_{1},s_{2}\}$, and $V(r)=\{s_{0},s_{2}\}$.
At $s_{0}$, the following individual knowledge about $a$, $b$, and $c$ are satisfied:
  • $K_{a}(p\to q),\neg K_{a}(q\to r),\neg K_{a}(p\to r)$;
  • $\neg K_{b}(p\to q),K_{b}(q\to r),\neg K_{b}(p\to r)$;
  • $\neg K_{c}(p\to q),\neg K_{c}(q\to r),\neg K_{c}(p\to r)$.
During knowledge pooling, the flow of information from server $a$ to customer $c$ or vice versa is crucial for generating distinct agent-dependent knowledge outcomes. Agent $c$’s knowledge can be augmented based on the classification from $a$, while $a$’s knowledge dependent on $c$ remains the same:
  • $K^{c}_{a}(p\to q),\neg K^{c}_{a}(q\to r),\neg K^{c}_{a}(p\to r)$;
  • $K^{a}_{c}(p\to q),\neg K^{a}_{c}(q\to r),\neg K^{a}_{c}(p\to r)$.
The further information pooling from server $b$ to $c$ leads to:
  • $K_{c}^{ab}(p\to q),K_{c}^{ab}(q\to r),K_{c}^{ab}(p\to r)$ [^1] The truth condition of $K_{c}^{ab}$ can be roughly understood as: $M,w\models K_{c}^{ab}\text{ iff }(R_{c}\cap\equiv_{a:w}\cap\equiv_{b:w})[w]\subseteq||\varphi||.$
Customer $c$ now knows the full information the rules of user behaviors. However, it does not imply that the customer $c$ is supposed to know the dataset of the users:
  • $\neg K_{c}^{ab}p$.
In contrast, distributed knowledge generated through information pooling does not merely copy the individual agent-based insights; it produces knowledge that extends well beyond the possible knowable limit of all individuals’ understanding.
  • $D_{\{a,b,c\}}(p\to q),D_{\{a,b,c\}}(q\to r),D_{\{a,b,c\}}(p\to r)$;
  • $D_{\{a,b,c\}}p,D_{\{a,b,c\}}q,D_{\{a,b,c\}}r$.
$s_{0}$$p,q,r$$s_{1}$$p,q,\neg r$$s_{2}$$\neg p,q,r$$s_{3}$$p,\neg q,\neg r$$s_{4}$$\neg p,\neg q,\neg r$$ac$$ac$$bc$$bc$
Figure 1: Knowledge sharing: Beyond the aggregation of arbitrary information for distributed knowledge.
Essentially, agents only share what they have known. For example, agent $a$ shares her knowledge of $p\to q$ with $c$, which $c$ does not know. They cautiously avoid extending to what they are uncertain. Neither $a$ nor $c$ can verify or falsify $p$ nor $q$; this information is still uncertain even after they share knowledge with each other. Distributed knowledge represents a special boundary of information pooling, characterized by the aggregation of all conceivable information into collective knowledge, even if gathering those uncertain goes beyond what can be verified. Conversely, knowledge pooling is specifically aimed at gathering confirmed or known information between agents.
A primary concern in knowledge pooling is about communication. The expression $K_{a}^{b}(p\wedge\neg K_{a}p)$, a variant of Moore sentence [22, 21], remains logically consistent. However, the interpretation of “$p$ is true but agent $a$ does not know $p$” becomes problematic because it leads to unsuccessful knowledge sharing, especially after agents have fully communicated with each other [22, 16]. A potential solution to this challenge is to introduce a dynamic mechanism for knowledge sharing. After knowledge sharing, $K_{a}^{b}(p\wedge\neg K_{a}p)$ will become inconsistent, while $K_{a}^{b}(p\wedge K_{a}p)$, which indicates a successful delivery of knowledge, will becomes true. This effective communication of knowledge pooling is crucial in evaluating the ethical and normative implication of such interaction. These dynamic processes will be explored in the next Section 2.2 and their ethical considerations will be discussed in Section 3.
2.2 Knowledge Sharing: A New Type of Information Pooling
Information flows from one agent to another, called information pooling, is a dynamic process. It typically involves aggregating all available information from the participating agents through information intersection [1, 30, 13, 8], without necessarily confirming whether the information constitutes knowledge. To ensure the accuracy and integrity of communication, this section posits that information pooling ought to involve the sharing of knowledge. To express the concept of knowledge sharing between two agents, where agent $b$ receives agent $a$’s knowledge, we introduce the notation $a\smalltriangleright b$ to denote this dynamic communication. Accordingly, the symbol $[a\smalltriangleright b]$ represents the act of sender $a$ conveying its knowledge to receiver $b$. Now we add the formula $[a\smalltriangleright b]\varphi$ into the language and interpret this formula as follows:
Definition 2.4 (Updated Model).
Let $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},V\rangle$ be a model and $w\in W$ be a point in $M$. Then:
$M,w\models[a\smalltriangleright b]\varphi\textrm{ iff }M|_{a\smalltriangleright b},w\models\varphi$
where $M|_{a\smalltriangleright b}=\langle W,\{R^{*}_{i}\}_{i\in\mathcal{I}},V\rangle$ is a $(a\smalltriangleright b)$-updated model of knowledge sharing such that
[…]
    If $i\neq b$, then $R^{*}_{i}=R_{i}$;
[…]
    If $i=b$, then $(s,u)\in R_{b}^{*}$
[…]
      When $s,u\in R_{b}[w]$, $(s,u)\in R_{b}\cap\equiv_{a:w}$;
[…]
      When $s,u\not\in R_{b}[w]$, $(s,u)\in R_{b}$.
The primary feature of our updated mechanism is its directional nature. It only eliminates those epistemic relations accessible to the recipient that do not align with the sender’s knowledge. This approach assumes that the recipient’s knowledge is dependent on the knowledge shared by the sender.
In addition, this updating process is syntactic. The fundamental principle of this knowledge sharing update is represented by intersecting $b$’s knowledge $R_{b}$ with the equivalence set $\equiv_{a:w}$ which reflects the sender $a$’s knowledge locally at $w$. As a result, the model is modified according to syntactic expressions that represent the sender’s knowledge, ensuring that the receiver adjusts their knowledge while maintaining every thing else being equally characterising the sender’s knowledge, a variant of the concept known as “ceteris paribus” [27, 19, 18]. This maintains the integrity of the sender’s knowledge in the recipient’s understanding.
Example 2.5 (Sensitive Information, Cont’d).
Revisiting the model $M$ illustrated in Figure 1, we present some findings on Moore sentences that illuminate the differences between static and dynamic knowledge arising from knowledge sharing:
  • $K^{a}_{c}((p\to q)\wedge\neg K_{c}(p\to q))$ is true at $w$ in $M$;
  • $K^{a}_{c}((p\to q)\wedge\neg K_{c}(p\to q))$ is false at $w$ in $M|_{a\smalltriangleright b}$;
  • $K^{a}_{c}((p\to q)\wedge K_{c}(p\to q))$ is true at $w$ in $M|_{a\smalltriangleright b}$.
The update process of knowledge sharing implies two key points. Firstly, it categorizes the recipient’s knowledge based on the knowledge provided by the sender, as shown in Lemma 2.6.2. Secondly, the modification of agent-dependent knowledge does not influence the distributed knowledge, as demonstrated in Lemma 2.6.3.
Lemma 2.6.
Let $M|_{a\smalltriangleright b}=\langle W,\{R^{*}_{i}\}_{i\in\mathcal{I}},V\rangle$ be a $(a\smalltriangleright b)$-updated model.
[…]
    $M|_{a\smalltriangleright b}$ is still a model.
[…]
    $R_{a}\cap R_{b}\subseteq R^{*}_{b}$.
[…]
    $D_{G}=D_{G}^{*}$.
Proof 2.7.
[…]
    We only need to demonstrate that $R^{*}_{b}$ is still an equivalence relation. This requires $\equiv_{a}$ to be an equivalence relation, which is not difficult to be proven given Definition 2.4.
[…]
    We only need to show: $R_{a}\subseteq\equiv_{a}$. Let $uR_{a}s$ and $\varphi\in\mathcal{L}$.
    • Let $M,u\models K_{a}\varphi$. Due to the 4 axiom of $K_{a}$, we have $M,u\models K_{a}K_{a}\varphi$. Since $uR_{a}s$, so $M,s\models K_{a}\varphi$.
    • Let $M,u\not\models K_{a}\varphi$. Due to the 5 axiom of $K_{a}$, we have $M,u\models K_{a}\neg K_{a}\varphi$. Since $uR_{a}s$, so $M,s\models\neg K_{a}\varphi$, which is $M,s\not\models K_{a}\varphi$.
[…]
    We know $D_{G}^{*}=R^{*}_{b}\cap D_{G\setminus\{b\}}$. Due to $R^{*}_{b}=R_{b}\cap\equiv_{a}$, so $R^{*}_{b}\subseteq R_{b}$. Thus, $D_{G}^{*}\subseteq D_{G}$. On the other hand, due to Lemma 2.6.2, $R_{a}\cap R_{b}\subseteq R^{*}_{b}$, so $R_{a}\cap R_{b}\cap D_{G\setminus\{b\}}\subseteq R^{*}_{b}\cap D_{G\setminus\{b\}}$. This is $D_{G}\subseteq D^{*}_{G}$.
In this work, we adopt a non-reductionist approach [36] to characterise the properties of dynamic knowledge sharing. Traditionally, the reduction approach [33] has been the predominant method for modelling updates in knowledge. This typically involves a reduction axiom, specifically the axiom of dynamic iteration, which aims to simplify the iteration $[a\smalltriangleright b][c\smalltriangleright d]$ into a less complex format. However, such simplification is not feasible within our framework. Consequently, we explore an alternative non-reductionist approach.
Proposition 2.8.
These properties are valid in our updated model.
  • (Inv) $(p\to[a\smalltriangleright b]p)\wedge(\neg p\to[a\smalltriangleright b]\neg p)$
  • (Rev) $\neg[a\smalltriangleright b]\varphi\to[a\smalltriangleright b]\neg\varphi$
  • (D) $[a\smalltriangleright b]\neg\varphi\to\neg[a\smalltriangleright b]\varphi$
  • (Int+) $[a\smalltriangleright b]K_{b}\varphi\to\bigvee_{\psi\in\mathcal{L}}(K_{b}[a\smalltriangleright b](\psi\to\varphi)\wedge K_{a}[a\smalltriangleright b]\psi)$
  • (Int+) $K_{b}[a\smalltriangleright b]\varphi\to[a\smalltriangleright b]K_{b}\varphi$
  • (Int-) $[a\smalltriangleright b]K_{c}\varphi\leftrightarrow K_{c}[a\smalltriangleright b]\varphi$ where $b\neq c$
  • (K) $[a\smalltriangleright b](\varphi\to\psi)\to([a\smalltriangleright b]\varphi\to[a\smalltriangleright b]\psi)$
  • (Ns) From $\varphi$ infer $[a\smalltriangleright b]\varphi$
  • (Inc) From $\varphi\to[a\smalltriangleright b]\psi$ infer $K_{a}\varphi\to[a\smalltriangleright b]K_{b}\psi$
These properties establish a sound and complete logical system for dynamic knowledge sharing (detailed in Section 4.3). The Inv axiom ensures that truth assignments remain invariant in the models before and after an update. The Rev axiom reflects that the $(a\smalltriangleright b)$-sharing update always results in a unique model. The D axiom is crucial, illustrating the feasibility of $(a\smalltriangleright b)$-sharing. The Int+, Int+, Int- axioms, and the Inc rule illustrate the interactions between the knowledge and sharing. Finally, the K axiom and Ns rule are considered standard. These properties infer the following two standard results.
Proposition 2.9.
These properties are valid:
  • C: $[a\smalltriangleright b]\varphi\wedge[a\smalltriangleright b]\psi\to[a\smalltriangleright b](\varphi\wedge\psi)$
  • RK: From $\vdash\varphi\wedge\varphi^{\prime}\to\psi$, it infers $\vdash[a\smalltriangleright b]\varphi\wedge[a\smalltriangleright b]\varphi^{\prime}\to[a\smalltriangleright b]\psi$.
  • RM: From $\varphi_{1},\dots,\varphi_{n}\to[a\smalltriangleright b]\psi$ inferring $K_{c}\varphi_{1},\dots,K_{c}\varphi_{n}\to[a\smalltriangleright b]K_{c}\psi$.
Using a fragment of our language, we can illustrate how one can transfer and share one’s knowledge with others.
Proposition 2.10.
The following properties are valid for the fragment regarding $Prop^{+}$, the set of Boolean propositions that are only constructed from atomic propositions, negation, and conjunction.
  • (Boolean) $\varphi\leftrightarrow[a\smalltriangleright b]\varphi$ where $\varphi\in Prop^{+}$
  • (Remain) $K_{c}\varphi\to[a\smalltriangleright b]K_{c}\varphi$ where $\varphi\in Prop^{+}$
  • (Sharing) $K_{a}\varphi\to[a\smalltriangleright b]K_{b}\varphi$ where $\varphi\in Prop^{+}$
  • (Step) $[a\smalltriangleright b]K_{b}\varphi\to[a\smalltriangleright b][b\smalltriangleright c]K_{c}\varphi$ where $\varphi\in Prop^{+}$
  • (Dist) $[a\smalltriangleright b]K_{b}\varphi\to D_{\{a,b\}}[a\smalltriangleright b]\varphi$
Axioms Boolean, Remain, and Sharing explain the process through which the knowledge of agent $a$ is accurately conveyed and integrated into the knowledge of another agent. Thus, the knowledge pooling from agent $a$ to agent $b$ augments the knowledge of the recipient $b$, without similar influence on the sender $a$ nor any third party such as agent $c$. This represents a unidirectional flow of information. The Dist axiom further presents the unidirectional flow of knowledge pooling can be transferred through multi-agent interactions.
Axiom Dist indicates that the knowledge derived from the pooling process contributes to the collective understanding shared from the sender to the recipient. However, it is not implied that all distributed knowledge can be generated from the knowledge pooling process. This distinction is clarified in Example 2.12 presented below. Through this, knowledge pooling reaches a new type of collective knowledge that bridges individual knowledge and distributed knowledge.
In this multi-agent context, our goal is to illustrate that the resolution of knowledge within the knowledge pooling process is not merely a matter of aggregating information arbitrarily. Consider a model $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},V\rangle$. The information resolution operator $R^{i}_{G}\varphi$, which represents the pooling of arbitrary information, is defined as per existing literature [1]:
$M,w\models R^{i}_{G}\varphi\textrm{ iff }M|_{G},w\models\varphi,$
where $M|_{G}=\langle W,\{R_{a}|_{G}\}_{a\in\mathcal{I}},V\rangle$ such that
$R_{a}|_{G}=\begin{cases}\bigcap_{b\in G}R_{b}&\quad a\in G,\\ R_{a}&\quad\textrm{otherwise}.\end{cases}$
While, let $G=\{a_{1},\dots,a_{n}\}$, we can define the notion of knowledge resolution by our operator of collective knowledge pooling:
$R_{G}\varphi:=[a_{1}\smalltriangleright a_{2}]\dots[a_{n-1}\smalltriangleright a_{n}][a_{n}\smalltriangleright a_{n-1}]\dots[a_{2}\smalltriangleright a_{1}]\varphi.$
The following statement illustrates a crucial distinction from knowledge to information pooling:
  • (IntR) $R_{G}K_{a}\varphi\to R^{i}_{G}K_{a}\varphi$.
Example 2.12 shows that the converse of this axiom is not possible. Besides, we can define another operator of individual knowledge pooling:
$R^{a_{1}}_{G}\varphi:=[a_{1}\smalltriangleright a_{2}]\dots[a_{n-1}\smalltriangleright a_{n}]\varphi.$
Now we have the following relationship between individual knowledge, “everybody knows” [33], individual sharing, collective sharing, and information pooling.
Proposition 2.11.
These properties are sound:
[…]
    $K_{a_{1}}\varphi\to R^{a_{1}}_{G}E_{G}\varphi$ where $\varphi\in Prop^{+}$
[…]
    $R^{a_{1}}_{G}E_{G}\varphi\to R_{G}E_{G}\varphi$ where $\varphi\in Prop^{+}$
[…]
    $E_{G}\varphi\to R_{G}E_{G}\varphi$ where $\varphi\in Prop^{+}$
[…]
    $R_{G}E_{G}\varphi\to R^{i}_{G}E_{G}\varphi$ where $\varphi\in Prop^{+}$
Example 2.12.
The model $M$ in Figure 2 is a counter-example of $R^{i}_{G}E_{G}\varphi\to R_{G}E_{G}\varphi$. In this model, the information resolution pools all consistent but unknown information gathering from agent $a$ and $b$ to have $R_{a}\cap R_{b}=\{(s_{0},s_{0})\}$. Therefore,
  • $M,s_{0}\models R^{i}_{\{a,b\}}E_{\{a,b\}}(p\wedge q\wedge r)$.
However, as both agents possess the same individual knowledge – each knows $p$ but nothing else – this results in:
  • $M,s_{0}\models R_{\{a,b\}}E_{\{a,b\}}p$,
  • $M,s_{0}\not\models R_{\{a,b\}}E_{\{a,b\}}q$, and
  • $M,s_{0}\not\models R_{\{a,b\}}E_{\{a,b\}}r$.
$s_{0}$ $p,q,r$$s_{1}$ $p,\neg q,\neg r$$s_{2}$ $p,\neg q,r$$s_{3}$ $p,q,\neg r$$a$$b$$b$
Figure 2: From information pooling to knowledge sharing.
This Example 2.12 demonstrates that knowledge pooling results in a distinct form of collective knowledge, diverging from the distributed knowledge generated through information resolution. The relationship among individual knowledge, the knowledge achieved by knowledge pooling, and the knowledge obtained by information pooling can be outlined as follows: For $\varphi\in Prop^{+}$,
$E_{G}\varphi\to K_{a}\varphi\to R^{a}_{G}E_{G}\varphi\to R_{G}E_{G}\varphi\to R^{i}_{G}E_{G}\varphi\leftrightarrow D_{G}R^{i}_{G}\varphi.$
An important aspect of knowledge pooling is that, despite the various directions in which it can occur, there always exists a resolution that obtains all knowledge from the participating agents, as illustrated in Example 2.13 below.
Example 2.13 (Sensitive Information, Cont’d).
Starting with the initial model presented in Figure 1, knowledge sharing towards the customer $c$ can proceed in two directions: either server $a$ or server $b$ initiates sharing their knowledge. This leads to distinct updated models, as shown in Figure 3:
  • $M,s_{0}\models[a\smalltriangleright c]K_{c}(p\to q)$;
  • $M,s_{0}\models[b\smalltriangleright c]K_{c}(q\to r)$.
$s_{0}$$p,q,r$$s_{1}$$p,q,\neg r$$s_{2}$$\neg p,q,r$$s_{3}$$p,\neg q,\neg r$$s_{4}$$\neg p,\neg q,\neg r$$ac$$ac$$b$$b$
(a) The impact of $a$’s knowledge sharing on $c$’s knowledge.
$s_{0}$$p,q,r$$s_{1}$$p,q,\neg r$$s_{2}$$\neg p,q,r$$s_{3}$$p,\neg q,\neg r$$s_{4}$$\neg p,\neg q,\neg r$$a$$ac$$bc$$bc$
(b) The impact of $b$’s knowledge sharing on $c$’s knowledge.
Figure 3: Two different directions of knowledge sharing
Within the framework of knowledge resolution, the concept of “everybody knows” knowledge [33] is achievable:
  • $M,s_{0}\models R^{a}_{\{a,b,c\}}E_{\{a,b,c\}}(p\to q)$;
  • $M,s_{0}\not\models R^{a}_{\{a,b,c\}}E_{\{a,b,c\}}(q\to r)$;
  • $M,s_{0}\models R^{b}_{\{a,b,c\}}E_{\{a,b,c\}}(q\to r)$;
  • $M,s_{0}\not\models R^{b}_{\{a,b,c\}}E_{\{a,b,c\}}(p\to q)$;
  • $M,s_{0}\models R_{\{a,b,c\}}(E_{\{a,b,c\}}(p\to q)\wedge E_{\{a,b,c\}}(q\to r)\wedge E_{\{a,b,c\}}(p\to r))$.
The concept of knowledge resolution highlights a crucial insight from knowledge pooling: its capacity to facilitate a consensus among a group  [^2] If we incorporate the notion of common knowledge [33] into our framework, we can validate the following statement about our knowledge resolution: $R_{G}E_{G}\varphi\leftrightarrow R_{G}C_{G}\varphi,$ where $C_{G}$ is the modality of common knowledge in group $G$.. However, a knowledge resolution may not always be allowed, particularly in scenarios involving sensitive information or issues arising from epistemic implications. The next section will propose a formalization for permissible knowledge pooling to address the gap between epistemic and ethical implications.
3 Permissible Knowledge Pooling
3.1 Permission to Know: Static Norms
To capture ethical implications within our model, we incorporate the concept of ideal knowledge transition. In standard deontic logic, the Kanger reduction typically employs a propositional constant to identify states within the model as permissible or impermissible [17]. To examine the interplay between epistemic and ethical implications, we introduce a propositional constant that specifies which pairs of states in the model, referred to as knowledge transitions, are permissible. Those knowledge transitions considered permissible are thus classified as ideal knowledge transitions.
We expand our framework of knowledge pooling as described below. A structure $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},O,V\rangle$ is a deontic model when $M$ is a model in additional with a symmetric relation $O\subseteq W\times W$ with the property that $\emptyset\neq O\subseteq\bigcup_{a\in\mathcal{I}}R_{a}$. To represent an ideal transition, we introduce the constant ${\sf O}$ and extend our language, indicating the presence of a permissible way to know – suggesting that it is okay to gain knowledge in such a manner. This constant is interpreted at a point $w$ in a deontic model $M$ as follows.
$M,w\models{\sf O}\textrm{ iff }O[w]\neq\emptyset.$
This constant has this validity:
${\sf O}\to\hat{K_{a}}{\sf O}$
for some agent $a$ [^3] By incorporating the concept of common knowledge, the validity expressed by $C_{\mathcal{I}}\bigvee_{a\in\mathcal{I}}O$ captures the characteristics of $O\subseteq\bigcup_{a\in\mathcal{I}}R_{a}$ and $O\neq\emptyset$.
Example 3.1 (Sensitive Information, Cont’d).
Figure 1 illustrates the state pairs $(s_{1},s_{0})$ and $(s_{3},s_{0})$ along with their corresponding symmetric pairs, constitute the ideal relation $O$. These pairs are thus considered as permissible pathways of obtaining knowledge.
We introduce a static notion termed as weak permission, denoted by $P_{a}\varphi$, and define it as $K_{a}\varphi\wedge\hat{K_{a}}{\sf O}$. Thus, $P_{a}\varphi$ at a point $w$ in the deontic model $M$ is interpreted as:
$M,w\models P_{a}\varphi\textrm{ iff }R_{a}[w]\subseteq||\varphi||\textrm{ and }R_{a}[w]\cap O[w]\neq\emptyset$
This implies that when an agent $a$ is permitted to acquire knowledge, it means that there is at least one ideal way for her to access her knowledge. The formula $P_{a}\varphi$ is designed to captures the concept of having permission to know. It asserts that an agent $a$ is permitted to know $\varphi$ if and only if, $\varphi$ is part of agent $a$’s knowledge base, and there is at least one permissible way of knowing it. This concept of “permission to know” can be expressed using a static language.
Example 3.2 (Sensitive Information, Cont’d).
After server $a$ conveys her knowledge to customer $c$, it becomes permissible for the agent $c$ to know $p\to q$. Likewise, when server $b$ shares his knowledge with customer $c$, it becomes permissible for $c$ to know $q\to r$. These permissions are true at point $s_{0}$ in the models illustrated in Figure 1:
  • $M|_{a\smalltriangleright c},s_{0}\models P_{c}(p\to q)$;
  • $M|_{b\smalltriangleright c},s_{0}\models P_{c}(q\to r)$.
However, in the model in Figure (3a), the green and unique $c$-link will be deleted after the $(b\smalltriangleright c)$-update, so it is not permissible for $c$ to know $p\to r$.
  • For instance, $(M|_{a\smalltriangleright c})|_{b\smalltriangleright c},s_{0}\not\models P_{c}(p\to r)$.
The static concept of permission to know displays characteristics derived from the principles of free choice permission [20] and individual knowledge:
Proposition 3.3.
Here are some validities:
[…]
    RFC: $P_{a}\varphi\wedge P_{a}\psi\to P_{a}(\varphi\vee\psi)$
[…]
    MC: $P_{a}(\varphi\wedge\psi)\leftrightarrow P_{a}\varphi\wedge P_{a}\psi$
[…]
    K: $P_{a}(\varphi\to\psi)\to(P_{a}\varphi\to P_{a}\psi)$
[…]
    D: $\neg P_{a}\bot$
[…]
    T: $P_{a}\varphi\to\varphi$
[…]
    4: $P_{a}\varphi\to P_{a}P_{a}\varphi$
[…]
    RE: From $\varphi\to\psi$ it infers $P_{a}\varphi\to P_{a}\psi$
[…]
    NEC: From $\varphi$ it infers $P_{a}\varphi$
However, this concept avoids the infamous free choice permission paradox [20] and the 5 axiom:
Example 3.4.
Here are some invalidities for all $a\in\mathcal{I}$:
[…]
    FCP1: $P_{a}(\varphi\vee\psi)\to P_{a}\varphi\wedge P_{a}\psi$
[…]
    FCP2: $P_{a}\varphi\to P_{a}(\varphi\wedge\psi)$
[…]
    5: $\hat{P_{a}}\varphi\to P_{a}\hat{P_{a}}\varphi$
Significantly, the permission to access knowledge represents a novel bridge linking individual knowledge with distributed knowledge, as described below:
$K_{a}\varphi\to P_{a}\varphi\to D_{G}(\varphi\wedge{\sf O}).$
The dual concept of “permission to know” – “ought to know” can be defined in the usual manner: $O_{a}\varphi:=\hat{P_{a}}\varphi$. Thus, “ought to know” $O_{a}\varphi$ corresponds to a conditional $(K_{a}\neg\varphi\to K_{a}\neg{\sf O})$. This interpretation of “ought to know” avoids the infamous Åqvist’s paradox [2].
3.2 Permission to Knowledge Pooling: Dynamic Norms
Now we define a new type of permission – the permission for one to communicate their knowledge towards another. This introduces a novel dynamic dimension to the concept of permission, particularly in terms of communication, that is represented as $P(a\smalltriangleright b)$. It denotes that agent $a$ is permitted to share their knowledge with that of agent $b$. More precisely,
$P(a\smalltriangleright b)=_{df}[a\smalltriangleright b]\hat{K_{b}}{\sf O}$
Agent $a$ is allowed to transfer knowledge to agent $b$, if and only if the subsequent communication from $a$ to $b$ results in a permissible way for receiver $b$ to now.
Example 3.5 (Sensitive Information, Cont’d).
Within the use of the formulas like $P(a\smalltriangleright b)$, we can further explore the case of sensitive information in Example 3.2. In the model $M|_{a\smalltriangleright c}$, server $b$ still knows $q\to r$. When $b$ shares this knowledge with customer $c$, $c$ knows it as well. However, there is no legitimate means for her to having obtained this knowledge, as shown in Example 3.2:
  • $M|_{a\smalltriangleright c},s_{0}\not\models[b\smalltriangleright c]\hat{K_{c}}{\sf O}$.
From these observations, we deduce that the transfer of knowledge from $b$ to $c$ is impermissible.
  • $M|_{a\smalltriangleright c},s_{0}\not\models P(b\smalltriangleright c)$.
Knowledge pooling shares significant ties with the concept of “permissions to share knowledge”, particularly within multi-agent interaction scenarios.
Proposition 3.6.
Here are two validities:
[…]
    $[a\smalltriangleright b]P(b\smalltriangleright c)\to P(a\smalltriangleright c)$
[…]
    If $\vdash K_{a}\varphi\leftrightarrow K_{b}\varphi$ then $P(a\smalltriangleright c)\leftrightarrow P(b\smalltriangleright c)$
Proposition 3.6.1 demonstrates that, if it is permissible to share knowledge from $b$ to $c$ after $a$’s sharing towards $b$, a knowledge sharing from $a$ directly to $c$ is also permissible. Proposition 3.6.2 addresses the condition for replacing participants in permissible pooling: if the originating knowledge of the senders is identical, then the permissibility status of their knowledge pooling is the same. Therefore, Proposition 3.6.2 explains one scenario in which epistemic implications can affect ethical considerations. Example 3.7 below shows that substituting recipients with identical knowledge does not guarantee the preservation of the original permissibility status.
Example 3.7.
Here are some invalidities:
[…]
    If $K_{b}\varphi\leftrightarrow K_{c}\varphi$ then $P(a\smalltriangleright b)\leftrightarrow P(a\smalltriangleright c)$.
4 Expressivity, Axiomatization, and Completeness
This section presents several technical results, including the expressivity, soundness, and completenss of various axiomatizations of knowledge sharing.
4.1 Expressivity
Now we list different langauges of knowledge sharing as follows:
$\mathcal{L}_{IK}::\,$ $\varphi:=p\mid\varphi\wedge\varphi\mid K_{a}\varphi$
$\mathcal{L}_{AK}::\,$ $\varphi:=p\mid\varphi\wedge\varphi\mid K_{a}\varphi\mid K^{a}_{b}\varphi$
$\mathcal{L}_{DK}::\,$ $\varphi:=p\mid\varphi\wedge\varphi\mid K_{a}\varphi\mid K^{a}_{b}\varphi\mid D_{G}\varphi$
$\mathcal{L}_{IKS}::\,$ $\varphi:=p\mid\varphi\wedge\varphi\mid K_{a}\varphi\mid[a\smalltriangleright b]\varphi$
$\mathcal{L}_{AKS}::\,$ $\varphi:=p\mid\varphi\wedge\varphi\mid K_{a}\varphi\mid K^{a}_{b}\varphi\mid[a\smalltriangleright b]\varphi$
$\mathcal{L}_{DKS}::\,$ $\varphi:=p\mid\varphi\wedge\varphi\mid K_{a}\varphi\mid K^{a}_{b}\varphi\mid D_{G}\varphi\mid[a\smalltriangleright b]\varphi$
$\mathcal{L}_{DKR}::\,$ $\varphi:=p\mid\varphi\wedge\varphi\mid K_{a}\varphi\mid K^{a}_{b}\varphi\mid D_{G}\varphi\mid R^{i}_{G}\varphi\mid[a\smalltriangleright b]\varphi$
where $p$ is an element of the (countable) set $Prop$ of atomic propositions, $a,b\in\mathcal{I}$ are elements of the (finite) set $\mathcal{I}$ of agents, and $G\subseteq\mathcal{I}$ is a subset of agents.
Theorem 4.1.
The diagram in Figure 4 illustrates the expressive power, indicating that any language pointed to by an arrow is strictly more expressive than the language from which the arrow originates.
[…]
Figure 4: Expressivity
4.2 System AK of Agent-dependent Knowledge
The logical system AK of agent-dependent knowledge is presented in Figure 5. It includes two interaction axioms Int and Cl to connect between individual and agent-dependent knowledge.
(S5) The S5 proof system for individual knowledge [33] (AK) S5 axioms for agent-dependent knowledge (Int) $K_{a}\varphi\to K^{b}_{a}\varphi$ (Cl) $K^{b}_{a}\varphi\to\bigvee_{\psi\in\mathcal{L}}(K_{a}\psi\wedge K_{b}(\psi\to\varphi))$ (NECA) From $\varphi$ inferring $K^{b}_{a}\varphi$
Figure 5: Axiomatization AK
Theorem 4.2.
The system AK is sound and (strongly) complete.
Proposition 4.3.
The axiom Cl is valid.
Proof 4.4.
Let $M,w\models K^{b}_{a}\varphi$. This is $(R_{a}\cap\equiv_{b:w})[w]\subseteq||\varphi||$. Now we need to find a $\psi\in\mathcal{L}_{AK}$ to satisfy $R_{a}[w]\subseteq||\psi\to\varphi||$ and $R_{b}[w]\subseteq||\psi||$. Notice that $R_{b}[w]\subseteq\equiv_{b:w}[w]$ by Lemma 2.6.2 and $R_{b}[w]\neq\emptyset$. So we are able to find $\psi\in\mathcal{L}_{AK}$ such that $R_{b}[w]\subseteq||\psi||\subseteq\equiv_{b:w}[w]$. Now we need to show $R_{a}[w]\subseteq||\psi\to\varphi||$. If not, there is a $u\in R_{a}[w]$ such that $u\in||\psi\wedge\neg\varphi||$. However, from $u\in||\psi||\subseteq\equiv_{b:w}[w]$ and $u\in R_{a}[w]$, it infers $u\in||\varphi||$. This leads to a contradiction.
Definition 4.5 (Maximally Consistent Sets).
Given the system AK, for any $\Sigma\subseteq\mathcal{L}_{AK}$:
  • $\Sigma$ is AK-consistent iff each finite $X\subseteq\Sigma$ such that ⊬$\not\vdash_{\bf AK}\neg\bigwedge X$; otherwise $\Sigma$ is AK-inconsistent.
  • $\Sigma$ is a maximally AK-consistent subset iff it is AK-consistent and if $\Sigma\subset\Sigma^{\prime}\subseteq\mathcal{L}_{AK}$ then $\Sigma^{\prime}$ is AK-inconsistent.
We define $MCS$ be the set of all maximally AK-consistent subsets.
Lemma 4.6.
Let $\Sigma$ be a member of $MCS$. For any $\varphi,\psi\in\mathcal{L}_{AK}$:
  • $\varphi\not\in\Sigma\Leftrightarrow\neg\varphi\in\Sigma$.
  • $\varphi\wedge\psi\in\Sigma\Leftrightarrow\varphi,\psi\in\Sigma$.
  • $\Sigma$ is closed under the logic AK, such that
    $\varphi\in\Sigma$ and $\vdash_{\bf AK}\varphi\to\psi$ imply $\psi\in\Sigma$.
Definition 4.7 (Canonical Model).
Given the system AK, we construct the canonical model $M^{C}=\langle W^{C},\{R^{C}_{a}\}_{a\in\mathcal{I}},V^{C}\rangle$ as follows:
  • $W^{C}=MCS$ is all maximally AK-consistent subsets.
  • $R^{C}_{a}=\{(\Delta,\Theta)\subseteq W^{C}\times W^{C}\mid K_{a}\varphi\in\Delta\Rightarrow\varphi\in\Theta\}$
  • $V^{C}(p)=\{\Delta\in W^{C}\mid p\in\Delta\}$
Hence, $\equiv_{b:\Gamma}^{C}$ is defined as $\{(\Delta,\Theta)\subseteq W^{C}\times W^{C}\mid\forall\varphi\in K_{b}\Gamma(\varphi\in\Delta\Leftrightarrow\varphi\in\Theta)\}$. Concerns about the possibility of an empty intersection [13] between $R^{C}_{a}\cap\equiv_{b:\Gamma}^{C}$ do not arise in our canonical model.
Lemma 4.8 (Truth Lemma).
Let $M^{C}=\langle W^{C},\{R^{C}_{a}\}_{a\in\mathcal{I}},V^{C}\rangle$ be the canonical model with respect to logic AK. For any $\varphi\in\mathcal{L}_{AK}$ and $\Delta\in W^{C}$:
$M^{C},\Delta\models\varphi\text{ iff }\varphi\in\Delta.$
Proof 4.9.
We only focus on the case of agent-dependent knowledge $K^{b}_{a}\varphi$. The other cases are standard.
  • Suppose $K^{b}_{a}\varphi\in\Delta$. We want to show $(R^{C}_{a}\cap\equiv^{C}_{b:\Delta})[\Delta]\subseteq||\varphi||_{M^{C}}$. Let $\Theta\in(R_{a}\cap\equiv_{b:\Delta})[\Delta]$. From $K^{b}_{a}\varphi\in\Delta$ and Cl axiom they infer $K_{a}\psi\in\Delta$ and $K_{b}(\psi\to\varphi)\in\Delta$ for some $\psi\in\mathcal{L}_{AK}$. From $(\Delta,\Theta)\in R^{C}_{a}$ it then infers $\psi\in\Theta$. From $(\Delta,\Theta)\in\equiv_{b:\Delta}$ it infers $\psi\to\varphi\in\Theta$. So $\varphi\in\Theta$. By inductive hypothesis, it is $\Theta\in||\varphi||_{M^{C}}$.
  • Suppose $\neg K^{b}_{a}\varphi\in\Delta$. We need a $\Theta\in MCS$ such that $\Theta\in(R^{C}_{a}\cap\equiv^{C}_{b:\Delta})[\Delta]$ but $\Theta\not\in||\varphi||_{M^{C}}$. We construct such a $\Theta$ in the following way. We enumerate the formulas in the countable set $\mathcal{L}_{AK}$ as $\varphi_{1},\dots,\varphi_{n},\dots$ and construct $\Gamma_{i}$ for $i\in\mathbb{N}$ as follows:
    $\Gamma_{0}$ =$=\{\neg\varphi\}\cup\Delta_{0},$
    $\Gamma_{i+1}$ =$=\begin{cases}\Gamma_{i}\cup\{\varphi_{i}\mid K_{a}^{b}\varphi_{i}\in\Delta\}&\text{if this set is {\bf AK}-consistent; }\\ \Gamma_{i}&\text{else. }\end{cases}$
    where $\Delta_{0}=\{\psi\in\mathcal{L}_{AK}\mid K_{b}\psi\in\Delta\}$.
[…]
      We need to show $\Gamma_{i}$ where $i\in\mathbb{N}$ is consistent.
[…]
        $\Delta_{0}$ must be consistent. We want to show $\Gamma_{0}$ is consistent. If not, then it must be some $\psi_{1},\dots,\psi_{n}\in\Delta_{0}$ such that $\vdash\bigwedge_{1\leq i\leq n}\psi_{i}\to\varphi$. By the RK rule of $K^{b}_{a}$ modality, it infers $\vdash\bigwedge_{1\leq i\leq n}K^{b}_{a}\psi_{i}\to K^{b}_{a}\varphi$. Due to $K^{b}_{a}\psi_{i}\in\Delta$ for any $1\leq i\leq n$ and $\Delta\in MCS$, it leads to $K^{b}_{a}\varphi\in\Delta$. This leads to a contradiction.
[…]
        Every $\Gamma_{i+1}=\Gamma_{i}\cup\{\neg\varphi\}$ should be consistent. Otherwise, there are $\psi_{1},\dots,\psi_{n}\in\Gamma_{i}$ such that $\vdash\neg\varphi\to\neg\bigwedge_{1\leq i\leq n}\psi_{i}$. That is $\vdash\bigwedge_{1\leq i\leq n}\psi_{i}\to\varphi$. By the rule RK of $K^{b}_{a}$ modality, it infers $\vdash\bigwedge_{1\leq i\leq n}K_{a}^{b}\psi_{i}\to K_{a}^{b}\varphi$. Notice $\neg K^{b}_{a}\varphi\in\Delta$. By the closure of $\vdash$ in $MCS$, it infers $\neg\bigwedge_{1\leq i\leq n}K_{a}^{b}\psi_{i}\in\Delta$. However, $K_{a}^{b}\psi_{i}\in\Delta$ for each $1\leq i\leq n$. It leads to a contradiction that $\Delta$ is consistent.
[…]
      We define $\Gamma=\bigcup_{i\in\mathbb{B}}\Gamma_{i}$. We need to show $\Gamma\in MCS$. To verify that $\Gamma$ is consistent, the proof is similar to the previous case. The next is to show $\Gamma$ is maximal. If not, then there is a $\psi$ such that $\Gamma\cup\{\psi\}\supset\Gamma$ is still consistent. In such a case, $\Gamma\cup\{\psi\}\subseteq\Gamma$, which leads to a contradiction.
[…]
      We want to show $\Gamma\in(R^{C}_{a}\cap\equiv^{C}_{b:\Delta})[\Delta]$, which is
[…]
        $\forall\varphi\in\mathcal{L}_{AK}\,(K_{a}\varphi\in\Delta\Rightarrow\varphi\in\Gamma)$, and
[…]
        $\forall\varphi\in K_{b}\Delta\,(\varphi\in\Delta\Leftrightarrow\varphi\in\Gamma)$
      From the Int axiom, given any $K_{a}\varphi\in\Delta$ it infers $K^{b}_{a}\varphi\in\Delta$. By the construction of $\Gamma$, it gives $\varphi\in\Gamma$. So the condition (3a) is verified. The condition (3b) can be verified by the construction of $\Gamma$. Given any $K_{b}\varphi\in\Delta$, it is the case that $\varphi\in K_{b}\Delta$. By the T axiom of $K_{b}$ modality it is always $\varphi\in\Delta$, while by the construction of $\Gamma$ it is the case that $\varphi\in\Gamma$. So the condition (3b) is always satisfied.
4.3 System IKS of Knowledge Sharing
The dynamic logic IKS of knowledge sharing is presented in Figure 6. Notably, this system adopts a non-reductionist approach [36] to model knowledge sharing.
(S5) The proof system of individual knowledge (Inv) $(p\to[a\smalltriangleright b]p)\wedge(\neg p\to[a\smalltriangleright b]\neg p)$ (Int+) $[a\smalltriangleright b]K_{b}\varphi\to\bigvee_{\psi\in\mathcal{L}}(K_{b}[a\smalltriangleright b](\psi\to\varphi)\wedge K_{a}[a\smalltriangleright b]\psi)$ (Int+) $K_{b}[a\smalltriangleright b]\varphi\to[a\smalltriangleright b]K_{b}\varphi$ (Int-) $[a\smalltriangleright b]K_{c}\varphi\leftrightarrow K_{c}[a\smalltriangleright b]\varphi$ where $b\neq c$ (Rev) $\neg[a\smalltriangleright b]\varphi\to[a\smalltriangleright b]\neg\varphi$ (D) $[a\smalltriangleright b]\neg\varphi\to\neg[a\smalltriangleright b]\varphi$ (K) $[a\smalltriangleright b](\varphi\to\psi)\to([a\smalltriangleright b]\varphi\to[a\smalltriangleright b]\psi)$ (Rep) $[a\smalltriangleright b]\varphi\leftrightarrow[a\smalltriangleright b][a\smalltriangleright b]\varphi$ (Ns) from $\varphi$ infer $[a\smalltriangleright b]\varphi$ (Inc) from $\varphi\to[a\smalltriangleright b]\psi$ infer $K_{a}\varphi\to[a\smalltriangleright b]K_{b}\psi$
Figure 6: Axiomatization IKS with non-reduction axioms
Theorem 4.10.
The system IKS is sound and (strongly) complete.
Proposition 4.11.
The axioms and rules in IKS are valid.
Proof 4.12.
Now we prove the validities of all axioms and rules for the $[a\smalltriangleright b]$ modality.
  • The validity of the Inv axiom can be verified by the definition of updated model.
  • Suppose $M,w\models[a\smalltriangleright b]K_{b}\varphi$. This is equal to $R_{b}^{a\smalltriangleright b}[w]\subseteq||\varphi||^{*}$. Now we need to find a $\psi\in\mathcal{L}_{IKS}$ such that $R_{b}[w]\subseteq||\psi\to\varphi||^{*}$ and $R_{a}[w]\subseteq||\psi||^{*}$. As $R_{a}[w]\subseteq\equiv_{a:w}[w]$ and $R_{a}[w]\neq\emptyset$, we are able to find out such a $\psi\neq\bot$ such that $R_{a}[w]\subseteq||\psi||^{*}\subseteq\equiv_{a:w}[w]$. Now we need to show $R_{b}[w]\subseteq||\psi\to\varphi||^{*}$. If not, there is a $u\in R^{b}[w]$ such that $u\in||\psi\wedge\neg\varphi||^{*}$. From $u\in||\psi||^{*}\subseteq\equiv_{a:w}[w]$ and $u\in R^{b}[w]$, it follows that $u\in R^{a\smalltriangleright b}_{b}\subseteq||\varphi||^{*}$. It then leads to a contradiction. So the Int+ axiom is valid.
  • This is straightforward by considering the definition of update.
  • Due to the definition of update, Int- is valid.
  • From $M,w\not\models[a\smalltriangleright b]\varphi$ it infers $M|_{a\smalltriangleright b},w\not\models\varphi$. This is $M|_{a\smalltriangleright b},w\models\neg\varphi$. So $M,w\models[a\smalltriangleright b]\neg\varphi$.
  • Suppose $M,w\models[a\smalltriangleright b]\bot$. So $M|_{a\smalltriangleright b},w\models\bot$, which is not possible.
  • Suppose $M|_{a\smalltriangleright b},w\models\varphi\to\psi$ and $M|_{a\smalltriangleright b},w\models\varphi$. So $M|_{a\smalltriangleright b},w\models\psi$. So the K axiom holds.
  • Suppose $M|_{a\smalltriangleright b},w\models\varphi$. Then, update the model $M|_{a\smalltriangleright b}$ by $a\smalltriangleright b$ does not change any truth assignments or binary relations in this model. So $(M|_{a\smalltriangleright b})|_{a\smalltriangleright b},w\models\varphi$.
  • Because $\varphi$ is valid. So, as in the standard proof [33], $\varphi$ is also valid in the updated models.
  • Suppose $||\varphi||\subseteq||\psi||^{a\smalltriangleright b}$. Let $R_{a}[w]\subseteq||\varphi||$ and $u\in R^{\smalltriangleright}_{b:c}[w]$. From $u\in R^{\smalltriangleright}_{b:c}[w]\subseteq\equiv_{a:w}$ it infers that $u\in||\psi||\Leftrightarrow w\in||\psi||$ for any $\psi\in K_{a}w$. From $R_{a}[w]\subseteq||\varphi||$ it leads to $u\in||\varphi||$. From the assumption $||\varphi||\subseteq||\psi||^{a\smalltriangleright b}$ it follows $u\in||\psi||^{a\smalltriangleright b}$. Now we have $R^{\smalltriangleright}_{b:c}[w]\subseteq||\psi||^{a\smalltriangleright b}$. So the Inc rule is valid.
Proposition 4.13.
These are valid in IKS:
  • C: $[a\smalltriangleright b]\varphi\wedge[a\smalltriangleright b]\psi\to[a\smalltriangleright b](\varphi\wedge\psi)$
  • RK: From $\vdash\varphi\wedge\varphi^{\prime}\to\psi$, it infers $\vdash[a\smalltriangleright b]\varphi\wedge[a\smalltriangleright b]\varphi^{\prime}\to[a\smalltriangleright b]\psi$.
  • RM: From $\varphi_{1},\dots,\varphi_{n}\to[a\smalltriangleright b]\psi$ inferring $K_{c}\varphi_{1},\dots,K_{c}\varphi_{n}\to[a\smalltriangleright b]K_{c}\psi$.
Proof 4.14.
[…]
Proposition 4.15.
These are derivable in IKS.
[…]
    (Boolean) $\varphi\leftrightarrow[a\smalltriangleright b]\varphi$ where $\varphi\in Prop^{+}$
[…]
    (Remain) $K_{c}\varphi\to[a\smalltriangleright b]K_{c}\varphi$ where $\varphi\in Prop^{+}$
[…]
    (Sharing) $K_{a}\varphi\to[a\smalltriangleright b]K_{b}\varphi$ where $\varphi\in Prop^{+}$
[…]
    (Step) $[a\smalltriangleright b]K_{b}\varphi\to[a\smalltriangleright b][b\smalltriangleright c]K_{c}\varphi$ where $\varphi\in Prop^{+}$
Proof 4.16.
[…]
    Boolean can be derived by Inv, D, K, and Ns.
[…]
    Remain can be verified by Boolean, and RM.
[…]
    Sharing can be verified by Boolean and Inc.
[…]
    Step can be verified by Sharing, K, Ns, and MP.
We adopt the detour canonical method [36] to prove the completenss of this dynamic logic IKS.
Definition 4.17 (Extended Model).
An extended model $M$ for IKS is a structure $\langle W,\{R_{a}\}_{a\in\mathcal{I}},\{R_{a\smalltriangleright b}\}_{a,b\in\mathcal{I}},V\rangle$ where:
  • $(\langle W,\{R_{a}\}_{a\in\mathcal{I}},V\rangle)$ is a model for IKS;
  • For any $a,b\in\mathcal{I}$, $R_{a\smalltriangleright b}$ is a relation over $W$.
We call $(\langle W,\{R_{a}\}_{a\in\mathcal{I}},V\rangle)$ the core of the extended model $M$ and denote it as $M^{-}$.
The truth conditions are defined as follows:
[…]
Definition 4.18 (Normal extended model).
An extended model $M=\langle W,\{R_{a}\}_{a\in\mathcal{I}},\{R_{a\smalltriangleright b}\}_{a,b\in\mathcal{I}},V\rangle$ for IKS is called normal, if and only if for any $w,u\in M$ and any $a,b,c\in\mathcal{I}$ such that $a\neq c$:
  • (U-Functionality) For any $w\in W$, $w$ has a unique $(a\smalltriangleright b)$-successor, that is $|R_{a\smalltriangleright b}[w]|=1$.
  • (U-Invariance) If $wR_{a\smalltriangleright b}u$ then $\forall p\in Prop$: $w\in V(p)\Leftrightarrow u\in V(p)$.
  • (U-Zig-) If $wR_{a}u$, $wR_{b\smalltriangleright c}w^{\prime}$, and $uR_{b\smalltriangleright c}u^{\prime}$, then $w^{\prime}R_{a}u^{\prime}$.
  • (U-Zig+) If $wR^{b}_{c}u$, $wR_{b\smalltriangleright c}w^{\prime}$, and $uR_{b\smalltriangleright c}u^{\prime}$, then $w^{\prime}R_{c}u^{\prime}$.
  • (U-Zag-) If $w^{\prime}R_{a}u^{\prime}$ and $wR_{b\smalltriangleright c}w^{\prime}$, then $\exists u$ such that $wR_{a}u$ and $uR_{b\smalltriangleright c}u^{\prime}$.
  • (U-Zag+) If $w^{\prime}R_{c}u^{\prime}$ and $wR_{b\smalltriangleright c}w^{\prime}$, then $\exists u$ such that $wR^{b}_{c}u$ and $uR_{b\smalltriangleright c}u^{\prime}$.
Definition 4.19 (Bisimulation).
A binary relation $Z$ is called a $({a\smalltriangleright b})$-bisimulation between two pointed model $M,w$ and $N,u$, denoted as $M,w{\overset{a\smalltriangleright b}{\leftrightarroweq}}N,u$, when $wZu$ the following conditions of $w^{\prime}Zu^{\prime}$ hold:
  • (Invariance) $p\in V^{M}(w)$ iff $p\in V^{N}(u)$
  • (Zig-) If $wR_{c}u$ for some $u\in M$ then there is a $u^{\prime}\in N$ such that $w^{\prime}R_{c}u^{\prime}$ and $uZu^{\prime}$, where $c\neq b$.
  • (Zig+) If $wR^{a}_{b}u$ for some $u\in M$ then there is a $u^{\prime}\in N$ such that $w^{\prime}R_{b}u^{\prime}$ and $uZu^{\prime}$.
  • (Zag-) If $w^{\prime}R_{c}u^{\prime}$ for some $u^{\prime}\in N$ then there is a $u\in M$ such that $wR_{c}u$ and $uZu^{\prime}$, where $c\neq b$.
  • (Zag+) If $w^{\prime}R_{b}u^{\prime}$ for some $u^{\prime}\in N$ then there is a $u\in M$ such that $wR^{a}_{b}u$ and $uZu^{\prime}$.
Lemma 4.20.
[…]
$M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},w^{\prime}.$
Proof 4.21.
We need to show that $\overset{a\smalltriangleright b}{\leftrightarroweq}$ is a bisimulation.
  • (Invariance) That $p\in V^{M^{-}|_{a\smalltriangleright b}}(w)\Leftrightarrow p\in V^{M^{-}}(w^{\prime})$ is ensured by U-Invariance.
  • (Zig-) Let $wR_{c}u$ where $c\neq b$. By U-Functionality, there is a $u^{\prime}$ such that $uR_{a\smalltriangleright b}u^{\prime}$. By U-Zig-, $w^{\prime}R_{c}u^{\prime}$. From $uR_{a\smalltriangleright b}u^{\prime}$ and inductive hypothesis, it follows $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$.
  • (Zig+) Let $wR^{a}_{b}u$. By U-Functionality, there is a $u^{\prime}$ such that $uR_{a\smalltriangleright b}u^{\prime}$. By U-Zig+, it infers $w^{\prime}R_{c}u^{\prime}$. From $uR_{a\smalltriangleright b}u^{\prime}$ and inductive hypothesis, it follows $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$.
  • (Zag-) Let $w^{\prime}R_{c}u^{\prime}$ where $c\neq b$. By U-Zag-, there is a $u\in M^{-}$ such that $wR_{c}u$ and $uR_{a\smalltriangleright b}u^{\prime}$. From $uR_{a\smalltriangleright b}u^{\prime}$ and inductive hypothesis, it follows $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$.
  • (Zag+) Let $w^{\prime}R_{b}u^{\prime}$. From U-Zag+, there is a $u$ such that $wR^{a}_{b}u$ and $uR_{a\smalltriangleright b}u^{\prime}$. From $uR_{a\smalltriangleright b}u^{\prime}$ and inductive hypothesis, it follows $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$.
Lemma 4.22.
When $M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},w^{\prime}$, it is the case that
$M^{-}|_{a\smalltriangleright b},w\models\varphi\Leftrightarrow M^{-},w^{\prime}\models\varphi.$
Proof 4.23.
We prove this lemma by induction on the complexity of $\varphi\in\mathcal{L}_{IKS}$.
  • When $\varphi\in Prop$. This can be verified by U-invariance.
  • The case of negation and conjunction can be verified by inductive hypothesis.
  • When $\varphi=K_{c}\psi$ such that $c\neq b$.
    • Assume that $M^{-}|_{a\smalltriangleright b},w\models K_{c}\psi$, it means $R_{c}[w]\subseteq||\psi||^{a\smalltriangleright b}$. Let $w^{\prime}R_{c}u^{\prime}$ for some $u^{\prime}$. By Zag-, there is a $u\in M^{-}|_{a\smalltriangleright b}$ such that $wR_{c}u$ and $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$. From $R_{c}[w]\subseteq||\psi||^{a\smalltriangleright b}$ it infers that $u\in||\psi||^{a\smalltriangleright b}$. By inductive hypothesis and $u\in||\psi||^{a\smalltriangleright b}$, $M^{-},u^{\prime}\models\psi$. Now we conclude that $M^{-},u^{\prime}\models K_{c}\psi$.
    • Assume that $M^{-},w^{\prime}\models K_{c}\psi$ and $wR_{c}u$. By Zig- and $wR_{c}u$, there is a $u^{\prime}$ such that $w^{\prime}R_{c}u^{\prime}$ and $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$. From $M^{-},u^{\prime}\models K_{c}\psi$ and $w^{\prime}R_{c}u^{\prime}$, it follows $M^{-},u^{\prime}\models\psi$. Again, by inductive hypothesis, we have $M^{-}|_{a\smalltriangleright b},u\models\psi$. So it leads to $M^{-}|_{a\smalltriangleright b},w\models K_{c}\psi$.
  • When $\varphi=K_{b}\psi$.
    • Assume that $M^{-}|_{a\smalltriangleright b},w\models K_{b}\psi$, it means $R^{a}_{b}[w]\subseteq||\psi||^{M^{-}|_{a\smalltriangleright b}}$. We need to show $M^{-},w^{\prime}\models K_{b}\psi$ where $wR_{a\smalltriangleright b}w^{\prime}$. Let $u^{\prime}\in M^{-}$ such that $u^{\prime}\in R_{b}[w^{\prime}]$. Given $M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},w^{\prime}$ and $u^{\prime}\in R_{b}[w^{\prime}]$, due to the Zag+, there is a $u\in M^{-}|_{a\smalltriangleright b}$ such that $wR^{a}_{b}u$ and $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$. From $wR^{a}_{b}u$ it leads to $u\in||\psi||^{a\smalltriangleright b}$. Now, by inductive hypothesis, it follows $M^{-}|_{a\smalltriangleright b},u\models\varphi\Leftrightarrow M^{-},u^{\prime}\models\varphi$. So $u\in||\psi||^{M^{-}}$, which is desired.
    • Assume that $M^{-},w^{\prime}\models K_{b}\psi$, which means $R_{b}[w^{\prime}]\subseteq||\psi||^{M^{-}}$. We want to show $R^{a}_{b}[w]\subseteq||\psi||^{M^{-}|_{a\smalltriangleright b}}$. Let $u\in R^{a}_{b}[w]$. Given $M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},w^{\prime}$ and $u\in R^{a}_{b}[w]$, due to Zig+, there is a $u^{\prime}\in M^{-}$ such that $w^{\prime}R_{b}u^{\prime}$ and $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$. From $w^{\prime}R_{b}u^{\prime}$ it follows $u^{\prime}\in||\psi||^{M^{-}}$. According to inductive hypothesis, we have $M^{-}|_{a\smalltriangleright b},u\models\varphi\Leftrightarrow M^{-},u^{\prime}\models\varphi$. Hence $u^{\prime}\in||\psi||^{M^{-}|_{a\smalltriangleright b}}$, which is desired.
  • When $\varphi=[c\smalltriangleright d]\psi$. We aim to $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d},w\overset{c\smalltriangleright d}{\leftrightarroweq}M^{-}|_{c\smalltriangleright d},w^{\prime}$.
[…]
      (Invariance) From the invariance of $M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},w^{\prime}$, it leads to $p\in V^{M^{-}|_{a\smalltriangleright b}}(w)$ if and only if $p\in V^{M^{-}}(w^{\prime})$. By the definition of $(c\smalltriangleright d)$-update, it infers $p\in V^{(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d}}(w)$ if and only if $p\in V^{M^{-}|_{c\smalltriangleright d}}(w^{\prime})$.
[…]
      (Zig-) Let $wR_{e}u$ for some $u\in(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d}$, where $e\neq d$. Because of the definition of update, the unchanged $e$-links, it therefore $wR_{e}u$ for $w,u\in M^{-}|_{a\smalltriangleright b}$. Due to the Zig- of $M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},w^{\prime}$, there is a $u^{\prime}$ such that $w^{\prime}R_{e}u^{\prime}$ and $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$. By the inductive hypothesis, we can conclude with $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d},u\overset{c\smalltriangleright d}{\leftrightarroweq}M^{-}|_{c\smalltriangleright d},u^{\prime}$.
[…]
      (Zig+) Let $wR^{c}_{d}u$ for some $u\in(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d}$. So $wR_{d}u$ where $u\in M^{-}|_{a\smalltriangleright b}$. If $d\neq b$, then, similar to the previous application of Zig-, there is a $u^{\prime}$ such that $w^{\prime}R_{d}u^{\prime}$ and $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d},u\overset{c\smalltriangleright d}{\leftrightarroweq}M^{-}|_{c\smalltriangleright d},u^{\prime}$. If $d=b$, then it must be $wR^{a}_{b}u$ as $w,u\in M^{-}|_{a\smalltriangleright b}$. By applying Zig+, similarly, it concludes that there is a $u^{\prime}$ such that $w^{\prime}R_{d}u^{\prime}$ and $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d},u\overset{c\smalltriangleright d}{\leftrightarroweq}M^{-}|_{c\smalltriangleright d},u^{\prime}$.
[…]
      (Zag-) Let $w^{\prime}R_{e}u^{\prime}$ for some $u^{\prime}\in(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d}$ where $e\neq d$. Again, due to the definition of update, the unchanged $e$-links, it infers $w^{\prime}R_{e}u^{\prime}$ where $w^{\prime},u^{\prime}\in M^{-}|_{a\smalltriangleright b}$. Because of the Zag- of $M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},w^{\prime}$, there is a $u\in M^{-}|_{a\smalltriangleright b}$ such that $wR_{e}u$ and $M^{-}|_{a\smalltriangleright b},u\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u^{\prime}$. Again, due to the inductive hypothesis, we can conclude with $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d},u\overset{c\smalltriangleright d}{\leftrightarroweq}M^{-}|_{c\smalltriangleright d},u^{\prime}$.
[…]
      (Zag+) Let $w^{\prime}R_{d}u^{\prime}$ for some $u^{\prime}\in(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d}$. Due to the definition of update, the removal of the $d$-links, it infers $w^{\prime}R_{d}u^{\prime}$ for some $u^{\prime}\in M^{-}|_{a\smalltriangleright b}$. When $d\neq b$, similar to the application of Zag- in the previous case, there is a $u$ such that $wR_{d}u$ in $M^{-}|_{a\smalltriangleright b}$ and $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d},u\overset{c\smalltriangleright d}{\leftrightarroweq}M^{-}|_{c\smalltriangleright d},u^{\prime}$. We need to show $w\equiv_{c:w}u$. From $w^{\prime}R_{d}u^{\prime}$ in $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d}$ it infers $w^{\prime}\equiv_{c:w^{\prime}}u^{\prime}$. From the invariance of the $K$-modality between $w$ and $w^{\prime}$ (the previous case) and the inductive hypothesis between $u$ and $u^{\prime}$, it then leads to $w\equiv_{c:w}u$. So we have $wR^{c}_{d}u$. When $d=b$, by Zag+ and inductive hypothesis, there is a $u$ such that $wR^{a}_{b}u$ in $M^{-}|_{a\smalltriangleright b}$ and $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d},u\overset{c\smalltriangleright d}{\leftrightarroweq}M^{-}|_{c\smalltriangleright d},u^{\prime}$. Similar to the case when $d\neq b$, it has $wR^{c}_{d}u$.
    By using $(M^{-}|_{a\smalltriangleright b})|_{c\smalltriangleright d},w\overset{c\smalltriangleright d}{\leftrightarroweq}M^{-}|_{c\smalltriangleright d},w^{\prime}$ and the inductive hypothesis, the case of $[c\smalltriangleright d]$ is easily verified.
Theorem 4.24.
For any $\varphi\in\mathcal{L}_{IKS}$ in IKS and any normal extended model $M$:
$M,s\models^{\smalltriangleright}\varphi\Leftrightarrow M^{-},s\models\varphi.$
Proof 4.25.
We only need to discuss the case of $[a\smalltriangleright b]\varphi$.
  • Let $M,w\models^{\smalltriangleright}[a\smalltriangleright b]\varphi$. So $\forall u\in R^{a}_{b}[w]:M,u\models^{\smalltriangleright}\varphi$. By U-Functionality, there is a $u\in R_{a\smalltriangleright b}[w]$. So $M,u\models^{\smalltriangleright}\varphi$. From inductive hypothesis, we know $M^{-},u\models\varphi$. From $u\in R_{a\smalltriangleright b}[w]$ and Lemma 4.20, we know that $M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u$. Since IKS-formulas are invariant under bisimulation, that is the Lemma 4.22, we have $M^{-}|_{a\smalltriangleright b},w\models\varphi\Leftrightarrow M^{-},u\models\varphi$. So we have $M^{-},w\models[a\smalltriangleright b]\varphi$, which is desired.
  • Suppose $M,w\not\models^{\smalltriangleright}[a\smalltriangleright b]\varphi$. So there is a $u\in R_{a\smalltriangleright b}[w]$ such that $M,u\not\models^{\smalltriangleright}\varphi$. By inductive hypothesis, $M^{-},u\not\models\varphi$. We want to show $M^{-},w\not\models[a\smalltriangleright b]\varphi$, which is $M^{-}|_{a\smalltriangleright b},w\not\models\varphi$. Due to $u\in R_{a\smalltriangleright b}[w]$ and Lemma 4.20, we have $M^{-}|_{a\smalltriangleright b},w\overset{a\smalltriangleright b}{\leftrightarroweq}M^{-},u$. So $M^{-}|_{a\smalltriangleright b},w\not\models\varphi$ from $M^{-},u\not\models\varphi$ and the invariance of bisimulation, the Lemma 4.22, again.
[…]
  • $\langle W^{\smalltriangleright},\{R_{a}^{\smalltriangleright}\}_{a\in\mathcal{I}},V^{\smalltriangleright}\rangle$ is a standard canonical model for IKS.
  • $R^{\smalltriangleright}_{a:b}=\{(\Gamma,\Delta)\subseteq W^{\smalltriangleright}\times W^{\smalltriangleright}\mid\forall\varphi\in\mathcal{L}_{IKS}:[a\smalltriangleright b]\varphi\in\Gamma\Rightarrow\varphi\in\Delta\}.$
So $(M^{\smalltriangleright})^{-}=M^{C}$. As in the standard canonical method [9, 36], it is obvious that $\Gamma R^{\smalltriangleright}_{a:b}\Delta$ if and only if $\varphi\in\Delta\Rightarrow\langle a\smalltriangleright b\rangle\varphi\in\Gamma$, where $\langle a\smalltriangleright b\rangle\varphi:=\neg[a\smalltriangleright b]\neg\varphi$.
Lemma 4.27 (Truth Lemma w.r.t. $\models^{\smalltriangleright}$).
For any $\varphi\in\mathcal{L}_{IKS}$:
$M^{\smalltriangleright},\Delta\models^{\smalltriangleright}\varphi\Leftrightarrow\varphi\in\Delta$
Proof 4.28.
All cases of $\varphi$ are standard, however, we will still present the proof of $\varphi=[a\smalltriangleright b]\psi$ here.
  • Suppose $[a\smalltriangleright b]\psi\in\Delta$. We want to show $M^{\smalltriangleright},\Delta\models^{\smalltriangleright}[a\smalltriangleright b]\psi$, which is $\forall\Gamma\in R^{\smalltriangleright}_{a:b}[\Delta]:M^{\smalltriangleright},\Gamma\models^{\smalltriangleright}\psi$. Let $\Gamma\in R^{\smalltriangleright}_{a:b}[\Delta]$. So $\psi\in\Gamma$ by the definition of $R^{\smalltriangleright}_{a:b}$ and $[a\smalltriangleright b]\psi\in\Delta$. By inductive hypothesis, $M^{\smalltriangleright},\Gamma\models^{\smalltriangleright}\psi$.
  • Suppose $\neg[a\smalltriangleright b]\psi\in\Delta$. We want to find a $\Gamma$ such that $\Gamma\in R^{\smalltriangleright}_{a:b}[\Delta]$ and $\Gamma\not\in||\psi||^{M^{\smalltriangleright}}$. We enumerate the formulas in the countable set $\mathcal{L}_{IKS}$ as $\varphi_{1},\dots,\varphi_{n},\dots$ and construct $\Gamma_{i}$ for $i\in\mathbb{N}$ as follows:
    $\Gamma_{0}$ =$=\{\neg\psi\},$
    $\Gamma_{i+1}$ =$=\begin{cases}\Gamma_{i}\cup\{\varphi_{i}\mid[a\smalltriangleright b]\varphi_{i}\in\Delta\}&\text{if this set is {\bf IKS}-consistent; }\\ \Gamma_{i}&\text{else. }\end{cases}$
[…]
      $\Gamma_{i}$ ($i\in\mathbb{N}$) is consistent. Otherwise, there are $\psi_{1},\dots,\psi_{n}\in\Gamma_{i}$ such that $\vdash\neg\psi\to\neg\bigwedge_{1\leq i\leq n}\psi_{i}$. This is $\vdash\bigwedge_{1\leq i\leq n}\psi_{i}\to\psi$. By applying the rule RK of $[a\smalltriangleright b]$ modality, it infers $\vdash\bigwedge_{1\leq i\leq n}[a\smalltriangleright b]\psi_{i}\to[a\smalltriangleright b]\psi$. From $\neg[a\smalltriangleright b]\psi\in\Delta$ it infers $\neg\bigwedge_{1\leq i\leq n}[a\smalltriangleright b]\psi_{i}\in\Delta$, which contradicts to $[a\smalltriangleright b]\psi\in\Delta$ and $\Delta$ is in $MCS$.
[…]
      We define $\Gamma$ as $\bigcup_{i\in\mathbb{N}}\Gamma_{i}$. So $\Gamma\in MCS$ as the standard proof.
[…]
      $\Gamma\in R^{\smalltriangleright}_{a:b}[\Delta]$ and $\neg\psi\in\Gamma$ can be guaranteed by the construction of $\Gamma$.
Proposition 4.29 (U-Functionality).
Given any $\Delta\in M^{\smalltriangleright}$, there is a unique $(a\smalltriangleright b)$-successor.
Proof 4.30.
There are two steps for this proof.
  • The existence of a $(a\smalltriangleright b)$-successor. We enumerate the fomulas of the counterble set $\mathcal{L}_{IKS}$ as $\varphi_{1},\dots,\varphi_{n},\dots$ and construct $\Gamma_{i}$ for $i\in\mathbb{N}$ as follows.
    $\Gamma_{0}$ =$=\{\psi\in\mathcal{L}_{IKS}\mid[a\smalltriangleright b]\psi\in\Delta\},$
    $\Gamma_{i+1}$ =$=\begin{cases}\Gamma_{i}\cup\{\varphi_{i}\in\mathcal{L}_{IKS}\}&\text{if this set is {\bf IKS}-consistent; }\\ \Gamma_{i}&\text{else. }\end{cases}$
    The set $\Gamma_{0}$ is consistent. If not, there are $[a\smalltriangleright b]\psi_{0},[a\smalltriangleright b]\psi_{1},\dots,[a\smalltriangleright b]\psi_{m}\in\Delta$ such that $\vdash\bigwedge_{1\leq i\leq m}\psi_{i}\to\neg\psi_{0}$. By the RK rule of $[a\smalltriangleright b]$, it leads to $\vdash\bigwedge_{1\leq i\leq m}[a\smalltriangleright b]\psi_{i}\to[a\smalltriangleright b]\neg\psi_{0}$. Notice that $[a\smalltriangleright b]\psi_{1},\dots,[a\smalltriangleright b]\psi_{m}\in\Delta$, it then infers $[a\smalltriangleright b]\neg\psi_{0}\in\Delta$. According to the D axiom, it then has $\neg[a\smalltriangleright b]\psi_{0}\in\Delta$, which leads to a contradiction to the assumption $[a\smalltriangleright b]\psi_{0}\in\Delta$.
  • If $\Gamma,\Gamma^{\prime}$ are both $\Delta$’s $(a\smalltriangleright b)$-successors, we need to show that they are the same set of formulas. If not, there is a $\varphi\in\Gamma$ and $\neg\varphi\in\Gamma^{\prime}$. From $\Gamma^{\prime}\in R^{\smalltriangleright}_{a:b}[\Delta]$, we have $\neg[a\smalltriangleright b]\varphi\in\Delta$. Because of the Rev axiom, we then have $[a\smalltriangleright b]\neg\varphi\in\Delta$. Taking together with $\Gamma\in R^{\smalltriangleright}_{a:b}[\Delta]$, it infers $\neg\varphi\in\Gamma$, which leads to a contradiction to $\varphi\in\Gamma$.
Proposition 4.31 (U-Invariance).
If $\Gamma\in R^{\smalltriangleright}_{a:b}[\Delta]$ then $\forall p\in Prop:p\in\Delta\Leftrightarrow p\in\Gamma$.
Proof 4.32.
This proposition is guaranteed by the definition of $R^{\smalltriangleright}_{a:b}$ and the Inv axiom.
Proposition 4.33 (U-Zig-).
If $\Delta R^{\smalltriangleright}_{a}\Gamma$, $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$, and $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$, then $\Delta^{\prime}R^{\smalltriangleright}_{a}\Gamma^{\prime}$.
Proof 4.34.
We need to show: $\forall\varphi\in\mathcal{L}_{IKS}:K_{a}\varphi\in\Delta^{\prime}\Rightarrow\varphi\in\Gamma^{\prime}$. Suppose $K_{a}\varphi\in\Delta^{\prime}$. So $\langle b\smalltriangleright c\rangle K_{a}\varphi\in\Delta$ from $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$. By the Rev axiom, we have $[b\smalltriangleright c]K_{a}\varphi\in\Delta$. Notice that $a\neq c$. By the Int- axiom it infers $K_{a}[b\smalltriangleright c]\varphi\in\Delta$. By applying $\Delta R^{\smalltriangleright}_{a}\Gamma$ and $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$, it has $\varphi\in\Gamma^{\prime}$.
Proposition 4.35 (U-Zig+).
If $\Delta R^{b\smalltriangleright}_{c}\Gamma$, $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$, and $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$, then $\Delta^{\prime}R^{\smalltriangleright}_{c}\Gamma^{\prime}$.
Proof 4.36.
We want to prove that $K_{c}\varphi\in\Delta^{\prime}\Rightarrow\varphi\in\Gamma^{\prime}$. Let $K_{c}\varphi\in\Delta^{\prime}$. So $\langle b\smalltriangleright c\rangle K_{c}\varphi\in\Delta$ from $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$. By the Rev axiom, we have $[b\smalltriangleright c]K_{c}\varphi\in\Delta$. By applying the Int+ axiom, there is a $\psi\in\mathcal{L}_{IKS}$ such that $K_{c}[b\smalltriangleright c](\psi\to\varphi)\in\Delta$ and $K_{b}[b\smalltriangleright c]\psi\in\Delta$. So, from $\Delta R^{b\smalltriangleright}_{c}\Gamma$ and $K_{c}[b\smalltriangleright c](\psi\to\varphi)\in\Delta$, it follows $[b\smalltriangleright c](\psi\to\varphi)\in\Gamma$. From $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$ it leads to $\psi\to\varphi\in\Gamma^{\prime}$. On the other hand, from $\Delta R^{b\smalltriangleright}_{c}\Gamma$ it indicates $\Delta\equiv_{b:\Delta}\Gamma$, which means that $\forall\varphi\in K_{b}\Delta:\varphi\in\Delta\Leftrightarrow\varphi\in\Gamma$. We already have $K_{b}[b\smalltriangleright c]\psi\in\Delta$. From the T axiom of the $K_{b}$ modality it is the case that $[b\smalltriangleright c]\psi\in\Delta$. So $[b\smalltriangleright c]\psi\in\Gamma$. From $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$ it infers $\psi\in\Gamma^{\prime}$. Now we conclude with $\varphi\in\Gamma^{\prime}$.
Proposition 4.37 (U-Zag-).
If $\Delta^{\prime}R^{\smalltriangleright}_{a}\Gamma^{\prime}$ and $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$, then there is a $\Gamma\in MCS$ such that $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$ and $\Delta R^{\smalltriangleright}_{a}\Gamma$.
Proof 4.38.
Notice that $a\neq c$. We enumerate $\varphi_{1},\dots,\varphi_{n},\dots$ in $\mathcal{L}_{IKS}$ and construct $\Gamma_{i}$ ($i\in\mathbb{N}$) as follows:
$\Gamma_{0}$ =$=\{\varphi\in\mathcal{L}_{IKS}\mid K_{a}\varphi\in\Delta\}\cup\{\langle b\smalltriangleright c\rangle\varphi\in\mathcal{L}_{IKS}\mid\varphi\in\Gamma^{\prime}\},$
$\Gamma_{i+1}$ =$=\begin{cases}\Gamma_{i}\cup\{\varphi_{i}\in\mathcal{L}_{IKS}\}&\text{if this set is {\bf IKS}-consistent; }\\ \Gamma_{i}&\text{else. }\end{cases}$
  • $\Gamma_{0}$ is consistent. If not, there is $K_{a}\varphi_{1},\dots,K_{a}\varphi_{n}\in\Delta$ and $\psi_{1},\dots,\psi_{m}\in\Gamma^{\prime}$ such that $\vdash\bigwedge_{1\leq i\leq n}\varphi_{i}\wedge\bigwedge_{1\leq j\leq m}\langle b\smalltriangleright c\rangle\psi_{j}\to\bot$. This means $\vdash\bigwedge_{1\leq i\leq n}\varphi_{i}\to\neg\bigwedge_{1\leq j\leq m}\langle b\smalltriangleright c\rangle\psi_{j}$, which is $\vdash\bigwedge_{1\leq i\leq n}\varphi_{i}\to\bigvee_{1\leq j\leq m}[b\smalltriangleright c]\neg\psi_{j}$. By using the RM rule, it is $\vdash\bigwedge_{1\leq i\leq n}K_{a}\varphi_{i}\to\bigvee_{1\leq j\leq m}[b\smalltriangleright c]K_{a}\neg\psi_{j}$. From $K_{a}\varphi_{i}\in\Delta$ for any $i$, it has $\bigwedge_{1\leq i\leq n}K_{a}\varphi_{i}\in\Delta$. So $\bigvee_{1\leq j\leq m}[b\smalltriangleright c]K_{a}\neg\psi_{j}\in\Delta$. Suppose $[b\smalltriangleright c]K_{a}\neg\psi_{k}\in\Delta$ where $\psi_{k}\in\Gamma^{\prime}$. From $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$, it follows $K_{a}\neg\psi_{k}\in\Delta^{\prime}$. Because $\Delta^{\prime}R^{\smalltriangleright}_{a}\Gamma^{\prime}$, it leads to $\neg\psi_{k}\in\Gamma^{\prime}$, which is a contradiction.
  • $\Gamma_{i}$ for each $i\in\mathbb{N}$ is consistent. The proof is similar to the case of $\Gamma_{0}$.
  • We define $\Gamma=\bigcup_{i\in\mathbb{N}}\Gamma_{i}$. So $\Gamma\in MCS$.
  • The construction of $\Gamma$ ensures that $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$ and $\Delta R^{\smalltriangleright}_{a}\Gamma$.
Lemma 4.39.
If $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$ then $K_{b}\varphi\in\Delta^{\prime}\Rightarrow[b\smalltriangleright c]K_{c}\varphi\in\Delta$.
Proof 4.40.
When $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$, it is $\forall\varphi\in\mathcal{L}_{IKS}:\varphi\in\Delta^{\prime}\Rightarrow[b\smalltriangleright c]\varphi\in\Delta$ by using the Rev axiom. From the left-to-right side of the Rep axiom, it leads to $\forall\varphi\in\mathcal{L}_{IKS}:\varphi\in\Delta^{\prime}\Rightarrow[b\smalltriangleright c]\varphi\in\Delta^{\prime}$. From $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$, it has $\varphi\to[b\smalltriangleright c]\varphi$ valid. So $K_{b}\varphi\to[b\smalltriangleright c]K_{c}\varphi$ valid from the Inc rule. If $K_{b}\varphi\in\Delta^{\prime}$ then $[b\smalltriangleright c]K_{c}\varphi\in\Delta^{\prime}$. From $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$ and the Rev axiom again, $[b\smalltriangleright c][b\smalltriangleright c]K_{c}\varphi\in\Delta$. From the right-to-left side of the Rep axiom, it leads to $[b\smalltriangleright c]K_{c}\varphi\in\Delta$.
Lemma 4.41.
If $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$ and $\Delta^{\prime}R^{\smalltriangleright}_{c}\Gamma^{\prime}$ then $\Delta^{\prime}\equiv_{b:\Delta^{\prime}}\Gamma^{\prime}$.
Proof 4.42.
We need to show: $\forall\varphi\in K_{b}\Delta^{\prime}:\varphi\in\Delta^{\prime}\Leftrightarrow\varphi\in\Gamma^{\prime}$. Given $\varphi\in K_{b}\Delta^{\prime}$:
  • If $\varphi\in\Delta^{\prime}$. From $\varphi\in K_{b}\Delta^{\prime}$, it has $K_{b}\varphi\in\Delta^{\prime}$. So $[b\smalltriangleright c]K_{c}\varphi\in\Delta$ from Lemma 4.39. From $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$ it infers $K_{c}\varphi\in\Delta^{\prime}$. From $\Delta^{\prime}R^{\smalltriangleright}_{c}\Gamma^{\prime}$ it leads to $\varphi\in\Gamma^{\prime}$.
  • If $\varphi\in\Gamma^{\prime}$. From $\varphi\in K_{b}\Delta^{\prime}$, it has $K_{b}\varphi\in\Delta^{\prime}$. From the T axiom of the $K_{b}$ modality, it infers $\varphi\in\Delta^{\prime}$, which is desired.
Proposition 4.43 (U-Zag+).
If $\Delta^{\prime}R^{\smalltriangleright}_{c}\Gamma^{\prime}$ and $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$, then there is a $\Gamma\in MCS$ such that $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$ and $\Delta R^{b\smalltriangleright}_{c}\Gamma$.
Proof 4.44.
We enumerate $\varphi_{1},\dots,\varphi_{n},\dots$ in $\mathcal{L}_{IKS}$ and construct $\Gamma_{i}$ ($i\in\mathbb{N}$) as follows:
$\Gamma^{0}$ =$=\{\varphi\mid K_{c}\varphi\in\Delta\}\cup\{\langle b\smalltriangleright c\rangle\varphi\mid\varphi\in\Gamma^{\prime}\},$
$\Gamma_{0}$ $=\Gamma^{0}\cup\{\varphi\mid K_{b}\varphi\in\Delta^{\prime}\}$
$\Gamma_{i+1}$ =$=\begin{cases}\Gamma_{i}\cup\{\varphi_{i}\in\mathcal{L}_{IKS}\}&\text{if this set is {\bf IKS}-consistent; }\\ \Gamma_{i}&\text{else. }\end{cases}$
Similar to the proof in Proposition 4.37, $\Gamma^{0}$ is consistent. We want to show: $\Gamma_{0}$ is consistent. Otherwise, $\exists K_{b}\chi_{1},\dots,K_{b}\chi_{k}\in\Delta$, $\exists K_{c}\varphi_{1},\dots,K_{c}\varphi_{n}\in\Delta$, and $\exists\psi_{1},\dots,\psi_{m}\in\Gamma^{\prime}$ such that
$\vdash\bigwedge_{1\leq l\leq k}\chi_{l}\to\neg(\bigwedge_{1\leq i\leq n}\varphi_{i}\wedge\bigwedge_{1\leq j\leq m}\langle b\smalltriangleright c\rangle\psi_{j}).$
By the NEC rule of the $K_{b}$ modality, it leads to
$\vdash K_{b}\bigwedge_{1\leq l\leq k}\chi_{l}\to K_{b}\neg(\bigwedge_{1\leq i\leq n}\varphi_{i}\wedge\bigwedge_{1\leq j\leq m}\langle b\smalltriangleright c\rangle\psi_{j}).$
From $K_{b}\chi_{1},\dots,K_{b}\chi_{k}\in\Delta$, it is $K_{b}\neg(\bigwedge_{1\leq i\leq n}\varphi_{i}\wedge\bigwedge_{1\leq j\leq m}\langle b\smalltriangleright c\rangle\psi_{j}).$ By applying either the Int+ or the Int- axioms, it infers
$\bigwedge_{1\leq i\leq n}K_{b}\varphi_{i}\to\bigvee_{1\leq j\leq m}[b\smalltriangleright c]K_{b}\neg\psi_{j}\in\Delta.$
From $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$ it follows
$\bigwedge_{1\leq i\leq n}K_{b}\varphi_{i}\in\Delta\Rightarrow\bigvee_{1\leq j\leq m}K_{b}\neg\psi_{j}\in\Delta^{\prime}.$
So, there is a $\psi_{k}\in\Gamma^{\prime}$ such that
$\bigwedge_{1\leq i\leq n}K_{b}\varphi_{i}\in\Delta\Rightarrow K_{b}\neg\psi_{k}\in\Delta^{\prime}.$
From $\Delta^{\prime}R^{\smalltriangleright}_{c}\Gamma^{\prime}$ and $\Delta R^{\smalltriangleright}_{b:c}\Delta^{\prime}$, and Lemma 4.41, we know that $\Delta^{\prime}\equiv_{b:\Delta^{\prime}}\Gamma^{\prime}$. So, when $K_{b}\neg\psi_{k}\in\Delta^{\prime}$, due to the T axiom of the $K_{b}$ modality, it is the case that
$\bigwedge_{1\leq i\leq n}K_{b}\varphi_{i}\in\Delta\Rightarrow\neg\psi_{k}\in\Gamma^{\prime}.$
Notice that $\psi_{k}\in\Gamma^{\prime}$, there is a $\varphi_{i}$ such that $\neg K_{b}\varphi_{i}\in\Delta$. However, it is in conflict with the assumption. So $\Gamma_{0}$ is consistent too. The consistency of all other $\Gamma_{i}$ can be verified easily.
We define $\Gamma=\bigcup_{i\in\mathbb{N}}\Gamma_{i}$. So $\Gamma\in MCS$. The construction of $\Gamma$ ensures that $\Gamma R^{\smalltriangleright}_{b:c}\Gamma^{\prime}$ and $\Delta R^{\smalltriangleright}_{c}\Gamma$. It also ensures $\Delta R^{\smalltriangleright}_{b}\Gamma$. From the proof of Lemma 2, we know that $R^{\smalltriangleright}_{b}\subseteq\equiv_{b:\Delta}$. So $\Delta\equiv_{b:\Delta}\Gamma$.
Proposition 4.45.
$M^{\smalltriangleright}$ is a normal extended model of $M^{C}$.
Proof 4.46.
This proposition can be verified by Proposition 4.29, Proposition 4.31, Proposition 4.33, Proposition 4.35, Proposition 4.37, and Proposition 4.43.
Lemma 4.47 (Truth Lemma).
For any $\varphi\in\mathcal{L}_{IKS}$ and a point $w\in M^{c}$:
$\varphi\in w\Leftrightarrow M^{C},w\models\varphi$
Proof 4.48.
This lemma can be verified by using Theorem 4.24 and Lemma 4.27 and by $(M^{\smalltriangleright})^{-}=M^{C}$ from Proposition 4.45.
4.4 Other Systems
According to the languages given in Section 4.1, we can continue axiomatizing the logical systems of knowledge sharing as follows. The system IK is the S5 proof system of individual knowledge [33]. The system AKS contains all axioms and rules both in AK and IKS, along with these two interaction axioms:
  • $[a\smalltriangleright b]K_{b}\varphi\leftrightarrow K^{a}_{b}[a\smalltriangleright b]\varphi$ and
  • $[a\smalltriangleright b]K^{a}_{b}\varphi\leftrightarrow K^{a}_{b}[a\smalltriangleright b]\varphi$.
The system DK contains all axioms and rules in the classical proof system for distributed knowledge [33] and in the system AK. The system DKS contains all axioms and rules in DK and in ASK together with this axiom:
  • $[a\smalltriangleright b]D_{G}\varphi\leftrightarrow D_{G}[a\smalltriangleright b]\varphi$.
Proposition 4.49.
$[a\smalltriangleright b]K_{b}\varphi\to D_{\{a,b\}}[a\smalltriangleright b]\varphi$ is derivable in DKS.
Proof 4.50.
  • It can be verified by Int+ axiom with the interaction axiom of $D_{G}$ modality.
The system DKR extends DKS with the axioms and rules in the system of resolved information [1] as well as this axiom:
  • (IntR) $R_{G}K_{a}\varphi\to R^{i}_{G}K_{a}\varphi$.
Proposition 4.51.
IntR is valid.
Proof 4.52.
  • Suppose $M,w\models R_{G}\varphi$. This indicates $R_{a}\cap\equiv_{G\setminus\{a\}}\subseteq||\varphi||^{*}$ where $a\in G$. Notice that $\bigcap_{i\in G}R_{i}\subseteq R_{a}\cap\equiv_{G\setminus\{a\}}$ from the proof of Lemma 2.6.2. So $\bigcap_{i\in G}R_{i}\subseteq||\varphi||^{*}$ together from applying Boolean.
Proposition 4.53.
These are derivable in DKR:
[…]
    $K_{a_{1}}\varphi\to R^{a_{1}}_{G}E_{G}\varphi$ where $\varphi\in Prop^{+}$
[…]
    $R^{a_{1}}_{G}E_{G}\varphi\to R_{G}E_{G}\varphi$ where $\varphi\in Prop^{+}$
[…]
    $E_{G}\varphi\to R_{G}E_{G}\varphi$ where $\varphi\in Prop^{+}$
[…]
    $R_{G}E_{G}\varphi\to R^{i}_{G}E_{G}\varphi$ where $\varphi\in Prop^{+}$
Proof 4.54.
[…]
    The following statements are derived from the axioms and rules of Sharing, Remain, Ns, K, and MP:
    $K_{a_{1}}\varphi\to[a_{1}\smalltriangleright a_{2}](K_{a_{1}}\varphi\wedge K_{a_{2}}\varphi)$
    by Sharing and Remain
    $[a_{1}\smalltriangleright a_{2}](K_{a_{1}}\varphi\wedge K_{a_{2}}\varphi)\to[a_{1}\smalltriangleright a_{2}][a_{2}\smalltriangleright a_{3}](K_{a_{1}}\varphi\wedge K_{a_{2}}\varphi\wedge K_{a_{3}}\varphi)$
    by Sharing, Remain, Ns, K, and MP
    $\cdots$
    $[a_{1}\smalltriangleright a_{2}]\cdots[a_{n-2}\smalltriangleright a_{n-1}]\bigwedge_{1\leq i\leq n-1}K_{a_{i}}\varphi\to[a_{1}\smalltriangleright a_{2}]\cdots[a_{n-1}\smalltriangleright a_{n}]\bigwedge_{1\leq i\leq n}K_{a_{i}}\varphi$
    So we have $K_{a_{1}}\varphi\to[a_{1}\smalltriangleright a_{2}]\cdots[a_{n-1}\smalltriangleright a_{n}]\bigwedge_{1\leq i\leq n}K_{a_{i}}\varphi$, which is the desired result $K_{a_{1}}\varphi\to R_{G}E_{G}\varphi$.
[…]
    The following statements are derivable from Proposition 2.11.1, Ns, K, and MP:
    $E_{G}\varphi\to R^{a_{n}}_{G}E_{G}\varphi$ by Proposition 2.11.1
    $[a_{n-1}\smalltriangleright a_{n}](E_{G}\varphi\to R^{a_{n}}_{G}E_{G}\varphi)$ by Ns
    $[a_{n-1}\smalltriangleright a_{n}]E_{G}\varphi\to[a_{n-1}\smalltriangleright a_{n}]R^{a_{n}}_{G}E_{G}\varphi$ by K and MP
    $\dots$ $\dots$
    $R^{a_{1}}_{G}E_{G}\varphi\to R_{G}E_{G}\varphi$ by Ns, K and MP
[…]
    Similar to Proposition 2.11.1, by Sharing, Remain, Ns, K, and MP, this formula is derivable.
[…]
    This can be derived from the IntR axiom.
So we have the following implication in DKR:
$E_{G}\varphi\to K_{a}\varphi\to R^{a}_{G}E_{G}\varphi\to R_{G}E_{G}\varphi\to R^{i}_{G}E_{G}\varphi.$
5 Related Work
5.1 Methodology of Link Deletion
Our model update, focused on $(a\smalltriangleright b)$-knowledge sharing, selectively eliminates all $b$-links within the model that do not align with agent $a$’s knowledge. In the field of model update and modification, deletion is widely recognized as a core strategy, as evidenced by various methodologies including public announcement [24], semantic updates [35], product updates [6], upgrades [28], and sabotage updates [25] among others. Technical surveys detailing the results of their complexities can referred to [33, 26, 3]. Link deletion, in particular, is essential for defining strategies in information pooling, where each method of update is designed for a specific function: Product updates [6, 33] are agent-oriented, upgrades [28] focus on formula-specific modifications, and sabotage methods [25, 26] target the removal of arbitrary links without regard to specific expressions  [^4] A recent proposal introduces a language-specific approach to sabotage updates [3]. Furthermore, an exploration of updates from a language-specific perspective to those unrelated to language in sabotage modifications was presented in [29]. . In the domain of information pooling, algebraic intersection methods that consolidate all potential information relative to specific agents [1, 30, 13, 8] stand out. Unlike previous approaches, our method remains agent-specific in the syntax in its deletion of links, but notably underscores the semantic role of knowledge formulas in the update process.
5.2 Social Dependence
Our knowledge pooling model brings to light a dimension of dependence within social interactions, a concept extensively examined in various formal theories. The notion of belief dependence, as introduced by Delgrande et al. [15], focuses on belief change and revision, thereby framing epistemic dependence. However, it is formalized within conditional logics employing a static language framework. On a different note, Boella et al. [10] have proposed a formal model designed to explain the power dynamics inherent in social dependence, rooted in an agent’s reliance on others to achieve their goals. This model of normative dependence utilises Castelfranchi’s methodology of dependence graphs [14], albeit without integrating a logical system.
In contrast, our framework for permissible knowledge pooling introduces a concept of dependence that not only underscores the dynamics of knowledge among individuals but also explores the capacity for maintaining the permissibility of information sharing between agents. This approach sheds light on the complex network of interpersonal dependencies, examined through both epistemic and normative lenses. Thus, this approach holds promise for future explorations into understanding social hierarchies and power dynamics.
5.3 Permission to Know
Our methodology for permissible knowledge pooling draws on the established logic of permission, as introduced by van der Meyden [31], which identifies permissibility across pairs of states concerning the authorization to acquire knowledge. Traditionally, most models have formalized the concept of permission to know by highlighting permissibility in relation to states. Moreover, inspired from Aucher et al. [4], several existing studies view permission to know as contingent upon permission to communicate—postulating that knowledge acquired through communication should not result in a breach of permissible states, as supported by the research of Balbiani et al. [5], Li et al.  [23], and Van Ditmarsch and Seban [32].
Our approach introduces a different perspective on the interplay between the permission to know and the permission to communicate, emphasising that the former precedes the latter. In essence, our framework suggests that we are granted the permission to know prior to receiving the authorization to communicate. This conceptualization of permissible knowledge pooling is distinctly defined through the lens of permission to know, offering a fresh viewpoint on its relationship with communicative permissions.
6 Conclusions
This study explores two key aspects of multi-agent communication: knowledge pooling and the permission to share knowledge. Unlike the conventional approach to information pooling, which aggregates arbitrary information to form collective knowledge, our framework selectively integrates only verified and known information during the pooling process. This approach introduces a novel perspective to information pooling, uncovering a unique form of collective knowledge that bridges the gap between individual and distributed knowledge.
Furthermore, our investigation offers a unique viewpoint on the interaction between epistemic and ethical considerations, successfully establishing a framework that explains how epistemic states together with our normative entitlement to knowledge influence our actions. By examining the relationship between the permission to know and the permission to share knowledge, we provide insightful observations on this interplay.
In our formal framework, we have investigated a range of validities and expressivities related to knowledge pooling, distributed knowledge, and the permission to know, emphasizing their roles within permissible knowledge pooling. We also explore their axiomatizations in addition with the soundness and completeness of systems AK for agent-dependent knowledge and IKS for dynamic knowledge sharing. The soundness and completeness of other systems are straightforward based on the results of AK and IKS. We leave the exploration of computational complexity and decidability to future technical research.
Our framework for permissible knowledge pooling introduces various challenges in the formalization of social reasoning within multi-agent interactions. For example, investigating the applicability of our concept of knowledge pooling – as a form of social dependence – to interpret social power is of significant interest. This includes examining the correlation between the extent of knowledge sharing and the structure of social hierarchies. Additionally, incorporating the concept of common knowledge into our framework presents an opportunity to explore deeper into the relationship between knowledge pooling and the achievement of social consensus.
Acknowledgement
This work is supported by the Fonds National de la Recherche of Luxembourg through the project Deontic Logic for Epistemic Rights (OPEN O20/14776480).